### v1

# VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems

Abstract.
Spatialized vibrotactile feedback systems deliver tactile information by placing multiple vibrotactile actuators on the body. As increasing numbers of actuators are required to adequately convey information in complicated applications, haptic designers find it difficult to create such systems due to limited scalability of existing toolkits. We propose VibraForge, an open-source vibrotactile toolkit that supports up to 128 vibrotactile actuators. Each actuator is encapsulated within a self-contained vibration unit and driven by its own microcontroller. By leveraging a chain-connection method, each unit receives independent vibration commands from a control unit, with fine-grained control over intensity and frequency. We also designed a GUI Editor to expedite the authoring of spatial vibrotactile patterns. Technical evaluations show that vibration units reliably reproduce audio waveforms with low-latency and high-bandwidth data communication. Case studies of phonemic tactile display, virtual reality fitness training, and drone teleoperation demonstrate the potential usage of VibraForge within different domains.
Haptics, Vibrotactile, Toolkits, Wearable Devices
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Haptic devicesccs: Human-centered computing User interface toolkits
Figure 1. VibraForge is a scalable, open-source vibrotactile toolkit that supports the creation of spatialized vibrotactile feedback systems. Leveraging a chain-connection method, one control unit in the system can independently control up to 128 vibration units. Applications of the toolkit are shown on the right.
[…]
An overview of the VibraForge toolkit, that has control and vibration units, uses a chain-connection topology, and supports the creation of multi-actuator devices.
1. Introduction
Tactile feedback conveys sensory information to the human body through the sense of touch (El Saddik et al., 2011). Vibrotactile feedback, which produces tactile sensations via actuators vibrating on the skin, has gained much popularity because it is compact, affordable, and accessible (Choi and Kuchenbecker, 2013). It has been widely adopted in various domains, such as virtual reality (VR) (Wang and et al., 2021; Kaul and Rohs, 2017; bHaptics, 2024), robotics (Pacchierotti and Prattichizzo, 2024; Xia et al., 2022; Aggravi et al., 2018), rehabilitation (Bark et al., 2014; Prabhu et al., 2020; Lindeman et al., 2006), and accessibility (Witteveen et al., 2015; Jung et al., 2022; Reed et al., 2018).
To provide adequate feedback to users during complex tasks, spatialized vibrotactile feedback systems have been proposed (Lindeman et al., 2004; Louison et al., 2017). These systems utilize actuator positions as an additional channel to enhance information transmission. For example, Wang et al. designed a facial mask with 36 eccentric rotating mass (ERM) actuators to study the perception of localization and orientation of objects in VR  (Wang and et al., 2021). The bHaptics TactSuit X40 (bHaptics, 2024) utilized 40 ERMs on a vest to create localized feedback for immersive VR gaming experiences. The bHaptics TactSuit was also used by Xia et al. to convey hydrodynamic flow intensity during submarine teleoperation (Xia et al., 2022). Jung et al. designed a wireless haptic interface with 36 ERMs to convey detailed fingertip touch sensations on the upper arm for amputees (Jung et al., 2022). Sanchez et al. created OpenVNAVI, a navigation aid system with 128 ERMs to guide visually impaired people (Sánchez and Bohne, 2015). A 64-actuator tactile jacket was also created by Lemmens et al. to enrich movie viewing experiences (Lemmens et al., 2009). Similarly, West et al. created the Body:Suit:Score system with 60 ERMs to enhance musical experiences (West et al., 2019).
Despite the great potential of spatialized vibrotactile feedback systems, previous implementations of such systems revealed scalability and expressivity limitations. Systems that depend on direct connections between actuators and microcontroller units (MCUs) exhibit limited scalability, as the number of actuators is intrinsically restricted by the available GPIOs or PWM channels on a MCU (Wang and et al., 2021; Kaul and Rohs, 2017; Reed et al., 2018). Expanding the number of actuators requires multiple MCUs, resulting in exponential increases in system complexity and communication overhead. For systems that use custom drivers and data bus protocols to drive actuators indirectly, the multi-layer connections support more actuators but compromise the vibration expressivity (LI et al., 2013; Lemmens et al., 2009; West et al., 2019). These systems cannot render complex waveforms such as audio signals due to bandwidth limits, thereby limiting the types of actuators that can be used to simple actuators like ERMs. Moreover, while various toolkits have been developed to assist vibrotactile design, they typically employ direct or multi-layer connections and face the same limitations (e.g., TECHTILE (Minamizawa et al., 2012), VITAKI (Martínez et al., 2014), and VHP (Dementyev et al., 2021)).
To overcome these limitations, this work proposes VibraForge, an open-source vibrotactile toolkit that leverages a scalable, modular design to enable the rapid prototyping of vibrotactile systems. VibraForge consists of control units and vibration units. Control units are responsible for receiving vibration control commands from a Bluetooth server and sending them to vibration units. Vibration units are self-contained modules with an actuator driven by their own microcontroller. Vibration units are connected to a control unit via a chain-connection method, where a single GPIO pin on a control unit can control up to sixteen vibration units via a custom UART protocol. To better support spatial pattern authoring, we also designed an accompanying GUI Editor that enables haptic designers to intuitively author multi-actuator patterns.
Technical evaluations of VibraForge showed that the toolkit has a high bandwidth (200 Hz), low latency (16 ms), and can reliably render complex audio signals in real time. Through three case studies, we demonstrate that VibraForge not only supports the efficient replication of existing research prototypes such as phonemic displays (Reed et al., 2018), but also improves user performance in various applications such as virtual reality fitness gaming and robot teleoperation.
The contributions of this work are:
  • an open-source hardware and software toolkit, VibraForge, that supports the design of spatialized vibrotactile feedback systems,
  • technical evaluations of VibraForge, including power consumption, bandwidth, latency and acceleration, etc., and
  • case studies demonstrating replicated research prototypes and novel applications enabled by VibraForge.
2. Related Work
Herein, we provide an overview of existing vibrotactile toolkits, highlighting their strengths and limitations. Then, we summarize previous spatialized vibrotactile feedback systems and discuss the benefits and drawbacks of their design strategies. Lastly, we summarize existing software for vibration pattern design.
2.1. Existing Vibrotactile Toolkits
Vibrotactile actuators such as eccentric rotation mass vibration motors (ERM), linear resonant actuators (LRA), voice coil actuators (VCA), or piezoelectric actuators (PIEZO) create mechanical vibrations via electromagnetic motions. Each actuator type and model has distinct voltage ratings, frequency responses, and vibration characteristics, making it challenging for designers to explore different configurations for wearable tactile devices. To assist the design process, various vibrotactile toolkits have been developed. These toolkits abstract low-level technical details and offer intuitive design interfaces, so researchers and designers can concentrate on creating vibration effects without being overwhelmed by the complexities of electronic, mechanical, and software design.
Vibrotactile prototyping toolkits can be categorized into those for education purposes and those for design purposes. Toolkits for education purposes often have simple compact widgets and limited output channels so they can teach novice users about the basics of haptic feedback and the design of simple vibration effects. For example, the TECHTILE toolkit was composed of a microphone, a signal amplifier, and two actuators (Minamizawa et al., 2012). Audio signals were captured by the microphone, amplified through the amplifier, and delivered to the actuators. The Stereohaptics toolkit used a similar principle, enabling users to create, record, and playback simple haptic waveforms, making haptic design accessible through familiar audio tools (Israr et al., 2016).
Toolkits for design purposes often have more sophisticated hardware and complex functionality that enable researchers and designers to explore different configurations of actuators and vibration patterns. For example, VITAKI was an ERM-based toolkit targeting virtual reality and video game applications (Martínez et al., 2014). It used shift registers and PWM drivers to control up to sixteen ERMs (i.e., the number of channels available on the PWM drivers). TactJam supported up to eight actuators and was intended for collaborative design (Wittchen et al., 2022). With Syntacts, designers could design vibration effects in a GUI Editor and then send them through an audio amplifier to actuators for testing (Pezent et al., 2020). The Syntacts audio amplifier, however, only supported eight channels and was too bulky to wear on the body. To enable the creation of wearable devices, the VHP toolkit used a custom designed PCB and flexible PCB connectors so devices could fit to different body shapes (Dementyev et al., 2021). The Nordic nRF52840 microcontrollers used in the toolkit had three PWM modules that generated 12 channels of PWM signals to drive the actuators.
While existing toolkits provide rich features to support vibrotactile design, they are limited by their scalability, particularly in the number of actuators they can support. This is because the actuators are directly controlled via GPIO pins or PWM channels on MCUs, with most offering a maximum of 16 channels. As vibrotactile feedback continues to be used for more complicated tasks and applications, these toolkits cannot meet the demands of emerging systems. Therefore, when designing VibraForge, we drew insights from previous spatialized vibrotactile feedback systems and propose a new connection method to overcome these scalability limitations.
2.2. Spatialized Vibrotactile System Design
Two connection methods have been often deployed When building spatialized vibrotactile systems (Figure 2). Some have tried maintaining a ”direct connection” to each actuator by adding more controllers to a system and thus increasing the number of PWM channels. For example, Wang et al. used PWM outputs from four Arduino Mega2560 boards to drive their facial display (Wang and et al., 2021). Similarly, Kaul et al. used four Arduino Nanos to drive 17 actuators mounted on their HapticHead display (Kaul and Rohs, 2017). Reed et al. used three custom-built audio amplifier boards to send individual waveforms to 24 actuators on a sleeve to deliver speech phonemic cues to deaf users (Reed et al., 2018). Although complex waveforms could be sent to the actuators via direct connections, communication overhead and system complexity increased significantly when the system scaled up.
Figure 2. Depiction of the direct connection (left) and multi-layer connection (middle) methods commonly used in spatialized vibrotactile systems and our proposed solution (right), which uses a chain connection method.
[…]
This figure illustrates three different connection methods for vibrotactile actuators: direct connection using PWM, multi-layer connection with SPI and custom drivers, and a chain connection solution using UART between vibration units for scalability.
Others have explored using ”multi-layer connections” that employ custom driver boards to interface between the microcontrollers and actuators (Sánchez and Bohne, 2015; LI et al., 2013; Lemmens et al., 2009; West et al., 2019). On the input side, these driver boards received vibration commands from a main controller board via data bus protocols such as I2C (Sánchez and Bohne, 2015) or SPI (Lemmens et al., 2009; West et al., 2019). On the output side, they converted vibration commands to PWM signals and drove four (Lemmens et al., 2009), six (West et al., 2019), eight (Sánchez and Bohne, 2015), or sixteen (LI et al., 2013) actuators. By leveraging a multi-layer design, these systems were more scalable than previous solutions, however, the indirect control afforded by the data bus protocols made it impossible to process complex waveforms. This limited the expressivity of the actuators so only simple actuators like ERMs were used in these systems.
There is thus a trade-off between system scalability and actuator expressivity: a system that supports fine-grained control of individual actuators is difficult to scale, and vice versa. In this work, we propose a chain-connection method (Figure 2), where actuators are encapsulated into self-contained vibration units and driven by their own PCB drivers. Chains of vibration units are connected to a central MCU. Each PCB driver receives commands from the central MCU and generates PWM signals for its own actuator, thus supporting scalability. The high bandwidth and low latency control of vibration characteristics enables actuators to render complex waveforms, thus enhancing expressivity
2.3. Software for Vibration Pattern Design
Vibration pattern design is crucial in spatialized vibrotactile feedback systems. For a comprehensive review of previous design software and their public availability, please see (Terenti and Vatavu, 2023). Here we discuss the common interfaces used to design single-actuator waveforms and multi-actuator patterns.
For single-actuator design, one common method is to use predefined waveforms. Designers start with standard waveforms such as sine, square, or triangle, multiply them to generate new waveforms, and modify them using envelopes (Pezent et al., 2020). While this method provides authoring flexibility, it requires that designers have a deep understanding of how waveforms can be composed through combinations of standard signals. This makes it difficult for novice designers to use. Others have created editors that enable the direct manipulation of waveforms. With these editors, designers have access to two parallel windows to edit vibration amplitude and frequency. For example, in the Macaron Editor, designers could directly move keyframe dots to modify a waveform, making the design process more intuitive (Schneider and MacLean, 2016). This method was later adopted by several industrial software tools, including Meta Haptics Studio (Meta Platforms, 2023a).
For multi-actuator authoring, the design focus shifts from waveform editing to pattern design. Designers compose vibration paths across multiple actuators on a canvas and the system converts the drawn paths to a series of vibration commands that are executed over time. For example, Schneider et al. utilized phantom sensation illusions to interpolate points between actuators to produce more continuous vibration sensations in Mango (Schneider et al., 2015).
VibraForge’s GUI Editor combines single-actuator waveform design with a multi-actuator canvas to support the efficient design for spatialized systems. It provides a library of standard waveforms and envelops for composing complex waveforms. It is also compatible with existing tools such as Syntacts (Pezent et al., 2020) and VIREO (Terenti and Vatavu, 2023) to support the reuse of workflows and prior knowledge. The canvas shows a virtual layout of actuators that replicates the physical chain connections so designers can use it as reference when designing spatial patterns.
3. VibraForge Design
As the discussion of prior work highlighted, there is a trade-off between actuator expressivity and system scalability. We aimed to mitigate this trade-off by using a novel chain-connection method with custom designed vibration units and control units. This section provides an overview of the VibraForge toolkit [^1] The project resource can be found at: https://anonymous.4open.science/r/VibraForge, including its design requirements, vibration units, control units, chain-connection method, signal segmentation algorithm, and vibration pattern GUI Editor.
3.1. Design Requirements
VibraForge’s design requirements were drawn from design challenges discussed in prior work on wearable spatial vibrotactile design or were identified by the authors based on prior prototyping experiences.
From spatial vibrotactile design, three design requirements were proposed:
  • DR1: Scalability. The toolkit should support large quantities of actuators. Designers should be able to use different types of actuators such as ERM, LRA, and VCA.
  • DR2: Fine-Grained Control. The toolkit should provide fine-grained control of individual actuator vibration characteristics such as intensity and frequency to support expressivity. Regardless of the number of actuators employed in the system, the control mechanism should have a high bandwidth and low latency.
  • DR3: Usability. The toolkit should provide an intuitive design interface for authoring vibrotactile spatial patterns. To leverage designers’ prior design knowledge, the toolkit should be compatible with existing vibrotactile waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a).
From wearable design, two design requirements were proposed:
  • DR4: Flexibility. The toolkit should enable designers to modify settings, reprogram behaviors, and physically rearrange components without difficulty. This requirement was derived from design challenges faced by West et al. when attempting to relocate or replace components that were sewn onto a garment (West et al., 2019).
  • DR5: Portability. Devices built with the toolkit should be portable and powered by batteries, communicate with external devices via wireless protocols, and use lightweight components to ensure a comfortable wearing experience.
3.2. Vibration Units
The vibrations units were designed as self-contained modules that could drive actuators by themselves. Each vibration unit consisted of a vibrotactile actuator, a custom designed PCB, and 3D printed components (Figure 3). We designed three versions of the vibration unit to support a range of actuators (DR1).
Figure 3. (Left) Exploded views of the small vibration unit for X-axis LRAs, including the cap, the PCB, the PCB mount, the enclosure, and the bottom ring. Z-axis LRAs shared most components except the enclosure. (Right) Exploded view of the large vibration unit for VCAs with a similar structure. Note that the bottom ring was placed underneath the garment to attach the unit.
[…]
An exploded view of the LRA and VCA vibration units, showing the actuator, PCB, and 3D printed parts.
3.2.1. Vibrotactile Actuators
An X-axis, square-shaped, 9.5 mm x 9.5 mm LRA manufactured by the Jinlong Machinery and Electronics Company (Model L959535) (Machinery and Co., 2022) was chosen for the vibration unit. Its resonant frequency was 170 Hz. When powered by 0.9 Vrms, it had a maximum current of 145 mArms and an acceleration of 0.65±0.12 Grms. The other two actuators were a Z-axis LRA manufactured by Jinlong (Model G0832) (Machinery and Co., 2022) and a VCA manufactured by PUI Audio, Inc. (Model HD-VA3222) (PUI Audio, 2023). The Z-axis LRA was cylindrical-shaped, with an 8 mm diameter and a height of 3.2 mm. Its resonant frequency was 235 Hz. When powered by 1.8 Vrms, it had a maximum current of 90 mArms and an acceleration of 0.50±0.08 Grms. The VCA was 32 mm x 22 mm and had a frequency response range from 80 - 500 Hz. When driven by a 133 Hz 1.5 Vrms signal, it had a maximum acceleration of 2.52 Gp-p.
These actuators were chosen because they cover a variety of vibration directions and intensities, thus satisfying the tactile perception requirements of different body parts (i.e., spatial acuity (Elsayed et al., 2020) and perceived intensity (Verrillo et al., 1969)). Since they all share the same PCB schematics and similar enclosure structures, in the following sections we only describe the vibration unit designed for the X-axis LRA. The other designs can be seen in Figure 3.
3.2.2. Printed Circuit Board Design
Due to bandwidth limitations, as the number of actuators increases it becomes difficult to drive them using real-time PWM signals sent from a central unit. Therefore, we allocated an MCU to each vibration unit to drive each actuator. Each vibration unit had its own driver PCB to receive UART commands from a central board and provide fine-grained control over vibration characteristics. The driver PCB had a PIC16F18313 MCU, a DRV8837 H-bridge motor driver, two 4-pin JST-SH headers, two pogo pins, and supportive components (Figure 4). The MCU was used to transmit commands and generate waveforms. The JST headers were used as input and output for the power and signal connections. The number of connector pins was minimized to four (i.e., actuator power, MCU power, ground, and data). The pogo pins were mounted on the bottom to ensure direct contact with the actuator pads.
Figure 4. The circuit schematics and PCB board designs of vibration units.
[…]
Circuit schematics of a unit driver that generated two opposite waveforms based on the received control signals to modulate the actuator.
When the MCU received a ”start” signal, it sent a waveform to the H-bridge to modulate the actuator. We utilized the PWM generator inside the MCU for fine-grained control over the actuators’ intensities, frequencies, and waveforms. Because there is no address system for each unit, designers only need to load their program into the MCU once. They can then plug-and-play each unit as desired. The communication protocol is further described in Section 3.4.
3.2.3. 3D Printed Enclosures
The 3D printed enclosures were designed to enclose each vibrotactile actuator and PCB board and provide a method to easily attach and detach each vibration unit to a garment (DR4). Similar to the size of a coin, the enclosure was compact and lightweight, enabling for a comfortable wearing experience (DR5).
The enclosure consisted of four parts: a cap, a PCB mount, an enclosure, and a bottom ring (Figure 3). The cap employed a screw-nut fastening system, which added top-down pressure to stabilize the PCB. The PCB mount offered protection for the PCB and aligned the pogo pins directly with the actuator pads. This improved the robustness of the vibration unit as no soldering was required during assembly. The enclosure served as a protective housing for all the parts. It had two holes for JST connectors, as well as two asymmetrically positioned vents for passive heat dissipation. The bottom ring was designed to be positioned beneath the garment to maintain direct contact with the skin without the need for any sewing (DR4).
The same design can be easily adapted to other vibration unit structures for other types of actuators. In Figure 3, other vibration units that share the same PCB schematics and slightly different enclosure designs are shown.
3.3. Control Unit
The control unit was designed to be fully portable (DR6). It consisted of a PCB extension board with an ESP32 microcontroller (MCU), two LiPo batteries, and a 3D-printed enclosure (Figure 5). Multiple versions of control units were designed to accommodate different projects, ranging from simple arm bands to more spatialized body displays. They varied in the (a) number of chains supported, (b) model of the ESP32 board, and (c) battery capacity.
Figure 5. The control unit, which contained an ESP32 MCU, an extension board, chain connectors, and LiPo batteries.
[…]
This figure shows an ESP32 board mounted on the extension board in the middle of the control unit, and the 8 chain connection ports extending out from the control unit.
We designed a large control unit that supported up to 8 chains of vibration units and 16 units per chain, for a total of 128 vibration units.The extension board drew power from two serial 3.7 V 8200 mAh LiPo batteries. It stabilized the voltage at 5 V through two power regulators: a linear regulator that powered all the MCUs and a buck regulator that powered all actuators in the system. These separate power supplies ensured that MCUs operated on stable voltages without interference from the actuators’ power consumption. The extension board had eight JST connectors. Each connector was connected to a GPIO pin on the ESP32 MCU and drove a chain of vibration units. The MCU was an Adafruit Feather ESP32 V2222https://learn.adafruit.com/adafruit-esp32-feather-v2. It was responsible for converting external Bluetooth commands to UART commands and sending them to the corresponding vibration unit.
Since the number of required vibration units varies across different tasks, we also designed a small control unit with fewer supported chains. This compact unit design contained two 3.7 V 500 mAh LiPo batteries and an Adafruit QT Py ESP32-S3  [^3] https://learn.adafruit.com/adafruit-qt-py-esp32-s3/pinouts. This MCU had four general purpose GPIO pins that could be used for chain connections.
3.4. Chain-Connection Method and Data Communication
VibraForge differs from existing toolkits due to its use of a chain-connection method (Figure 6). This method enables the toolkit to address over 100 vibration units using a single control unit (DR1) and support the simultaneous control over the vibration characteristics of multiple actuators (DR2).
Figure 6. The multiple chain-connection method used within VibraForge, depicting the connections required for a given chain.
[…]
This figure shows where the control unit, the power converters, and unit drivers are located to illustrate the multiple chain-connected method.
As explained in Section 3.3, each GPIO pin on the ESP32 MCU was routed to a JST connector on the extension board. Because vibration units shared the same JST connectors, they could be connected in chains and receive commands through the same GPIO pin. There were two options for the data transmission design. The first option was to use data bus protocols such as CAN or I2C, where all the vibration units would receive the same command simultaneously and use the command’s address to determine whether the command was intended for itself. While this option would allow commands to be transmitted with a minimal latency, the vibration units would need to be programmed with different addresses prior to use, which would be burdensome when working with multiple units. Instead, we decided to send commands through each chain using a serial protocol, where each unit would receive a command, process it, and resend it to the next unit. While this may increase the overall latency, it allowed each unit to operate with a ”virtual address”, enabling independent plug-and-play control without additional programming.
Figure 7. The two-byte message bits definition and an example of how vibration commands are transmitted along a chain using the ”virtual address” system.
[…]
The figure shows the two-byte protocol that was developed to control each vibration unit. In both bytes, a 9-bit UART with parity-bit check was used to ensure message integrity. The address and start/stop bit were in the first message, and operational commands including the duty cycle, frequency, and waveform, were in the second message.
Vibration commands were transmitted via a UART protocol at 115.2 kHz with parity checks (Figure 7). Each command consisted of two bytes: the first byte encoded the address and start/stop status of the target vibration unit, and the second byte detailed the vibration characteristics. Upon receiving the first byte, a vibration unit checked the address. A non-zero address indicated the command was for subsequent units, prompting the unit to decrement the address and forward the message. This process continued until the address reached zero and signified the targeted unit, which then acted based on the start/stop bit: ’1’ (START) awaited the second byte, while ’0’ (STOP) disabled its actuator. When a unit was ”signified”, it waited for the second-byte message, which contained details about vibration characteristics, including the 16 intensity levels (evenly distributed from 0% to 100% duty cycle) and 8 frequencies (i.e., 123, 145, 170, 200, 235, 275, 322, and 384 Hz). The frequency levels were chosen so that each level was 1.18 times higher than the previous level. This followed Weber’s factor for vibrotactile frequency just-noticeable differences (Pongrac, 2006).
3.5. Signal Segmentation Algorithm
To render complex waveforms such as audio signals on vibration units (DR2), we implemented a signal segmentation algorithm that converted high-frequency input signals to low-frequency serial commands (Figure 8).For instance, consider an input waveform of 44100 Hz audio signals that composed of 200 Hz and 5 Hz sine waves. High-frequency components (200 Hz) in the waveform were extracted using a short-time Fourier transform (STFT), which captured how the dominant frequency changed over time. Low-frequency components (5 Hz) were extracted using a Hilbert transform, which captured the envelope of the waveform. Both components were downsampled to 200 Hz and sent to vibration units. Since our frequency levels ranged from 123 Hz to 384 Hz, the segmentation threshold was set to 100 Hz. Frequency components above 100 Hz were rendered using frequency levels, while components below 100 Hz were rendered using intensity levels. The technical evaluations in Section 4 demonstrated that this algorithm could reliably render complex waveforms such as audio signals using serial command control.
Figure 8. The signal segmentation workflow, showing how low frequency and high frequency components were extracted from a complex waveform.
[…]
The left figure shows multiplication waveform of 200Hz and 5Hz sine waves. The middle one shows STFT spectrogram with the highest energy surrounds 200Hz. The right one shows the envelope of the waveform.
3.6. GUI Editor for Designing Spatialized Patterns
We designed a GUI Editor (Figure 9) to enable designers to create spatialized vibration patterns (DR3). This editor differed from previous editors in that it combined the waveform editing from single-actuator design software with the canvas design used for multi-actuator design. The key components of the GUI Editor were:
Figure 9. The layout of the GUI Editor, with its Waveform Library, Waveform Visualizer, Canvas, and Timeline.
[…]
The image shows a software interface with four main sections: The Waveform Library in the left panel contained a list of different waveforms available. The Waveform Visualizer at the top graphically depicted the selected waveform’s shape and behavior over time. The Canvas for Vibration Units in the center are where vibration units were visually arranged and connected. The Timeline at the bottom displayed the sequence and timing of waveform actions and events.
  • Waveform Library: The Waveform Library contained waveform templates such as oscillators and envelopes. Designers could use them to create a custom waveform and save it to the Waveform Library. In addition, designers could import waveforms in the JSON format from popular waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a) so they could reuse previous designs (DR3).
  • Waveform Visualizer: The Waveform Visualizer served two purposes. When designers were creating a custom waveform, they could drag and drop waveform templates from the Waveform Library onto the Waveform Visualizer and it would show the combined waveform. When designers were composing vibration patterns for individual vibration units, the Waveform Visualizer also show the composed waveform for the selected unit.
  • Canvas: The Canvas enabled designers to drag-and-drop vibration units onto the canvas so they could be connected into chains that mimicked the physical layout of the vibration units. In addition, a ”Create New Chain” function enabled designers to generate vibration units in batches and arrange them using grid layouts. When clicking on a unit, its corresponding waveform would show up in the Waveform Visualizer.
  • Timeline: The Timeline served as a high-level summary of the vibration patterns. The blue blocks showed the range of time when the vibration unit was activated. By moving the slider left and right, the vibration units that were activated at that moment would be highlighted in the Canvas so designers could understand the vibration patterns over time. By clicking the ”Play” button, the waveform would be converted into vibration commands with timestamps and sent to the control unit via Bluetooth.
In a typical spatial pattern design workflow, designers would begin by creating a new design and specify the number of chains and actuators to be used in their system. The GUI would then generate a default Canvas with the designated number of actuators. Designers would then freely adjust the layout to match their physical device. Then, designers would assign waveforms to each vibration unit and specify the start and end time. Each waveform could be designed from scratch using the waveform templates in the Waveform Library, or imported from other software. To review the spatialized patterns, designers would click on any vibration unit on the Canvas and the Waveform Visualizer would show the corresponding waveform. Alternatively, designers could see a high-level summary of how the vibration patterns of multiple units would change over time by moving the Timeline slider.
Additionally, we also provided a Python server with VibraForge APIs, so that vibration units can be directly activated and controlled via code (DR3). Later in Section 5, Python servers are integrated into external programming environments (e.g., Unity) to trigger real-time feedback for applications.
4. Technical Evaluation
In this section, we summarize various performance metrics of the VibraForge toolkit to demonstrate how it fulfills the aforementioned design requirements. All tests below used vibration units of X-axis LRAs and the small control unit with default level-7 intensity (50% duty cycle) and 170 Hz frequency, unless otherwise stated.
4.1. Latency
Having fine-grained control (DR2) over the actuators in a vibrotactile system means that the data transmission of the system must have low latency and high bandwidth. To determine if the latency of VibraForge was below the human perceivable latency threshold of 45 ms (Vogels, 2004), we measured the end-to-end latency of a PC to a control unit (i.e., the transmission latency of the Bluetooth Low Energy protocol (BLE)), and of a control unit to a vibration unit (i.e., the latency of the custom UART protocol).
To measure the PC to control unit latency, a BLE write GATT command was sent and immediately followed by a BLE read GATT command. A Python timer recorded the write command sending time and read command receiving time and computed the difference as the latency. We repeated this measurement 100 times and found that the average round-trip transmission latency was $26.67\pm 6.53$ ms. Therefore, the one-way transmission latency was approximately $14$ ms. For the control unit to vibration unit latency, repeated measurements using an oscilloscope showed that each command took $125$ us to be transmitted from one unit to the next unit on a chain. Therefore, latency on the longest chain with 16 actuators would accumulate to $2$ ms. Adding the two measurements together would thus result in an overall latency of approximately $16$ ms, which would be virtually imperceptible by humans (Vogels, 2004).
4.2. Bandwidth
The bandwidth of VibraForge was the number of commands that could be sent and processed per second by control units and vibration units. Bandwidth differs from latency in that asynchronous BLE commands could be continually sent without needing to wait for the previous command to be received (Figure 10). Therefore, bandwidth represented the refresh rate of the vibration commands for each actuator. Since BLE commands (millisecond level) would be processed much slower than UART commands (microsecond level), the bandwidth bottleneck of the system depended on the processing speed of BLE commands on the control unit.
Figure 10. Data transmission bandwidth breakdown.
[…]
This figure shows the breakdown of data transmission bandwidth and latency, highlighting BLE (Bluetooth Low Energy) transmission and UART communication. The BLE transmission segments are separated by 14 milliseconds, followed by 2 milliseconds of UART communication. The bandwidth is marked as 5 milliseconds or 200 Hz, representing the frequency of data transmission cycles. The diagram visually represents how BLE and UART transmissions alternate in time.
To measure VibraForge’s bandwidth, we wrote a Python testing program to repetitively send standard BLE commands without latency, and another program on the ESP32 (i.e., control unit) to receive the commands and print the receiving time to the serial port. The command processing time was computed as the difference between the receiving time for each command. We sent $1,000$ BLE commands with five UART commands in each of them, and computed an average processing time as $2.96\pm 1.20$ ms. Therefore, for the actual operation of the toolkit, we set the BLE command sending delay to be $5$ ms with some tolerance. This means that the control unit could simultaneously update five vibration units at $200$ Hz refresh rate, satisfying the signal segmentation algorithm requirements.
Although vibration units were not directly driven by continuous audio signals from the central MCU, the high bandwidth ensured that they could approximate continuous waveforms using discrete commands (DR2).
4.3. Power Consumption and Battery Life
Power consumption measurements would provide an estimation of how long wearable devices built using VibraForge would function before recharging was necessary (DR6). Power consumption was measured using a bench power supply with a real-time measurement of voltage, current, and power. When only the control unit was connected, it consumed $106$ mA current. When multiple vibration units were connected to the control unit, each vibration unit consumed $2.5$ mA when idle. When a vibration unit was activated, the actuator consumed $150$ mA. Therefore, for a vibrotactile arm display with one control unit and two full chains of vibration units ($16x2)$, the idle current would be $186$ mA and a $500$ mAh LiPo battery would last $2.69$ hours. Even in extreme situations such as multimedia entertainment, running two actuators consistently would consume $486$ mA, leading to $1.03$ hours of operating time. For a larger system with four chains of vibration units ($416x$ 4), the idle current would be $266$ mA, so a $8200$ mAh LiPo battery would last $30.83$ hours. In an extreme situation where eight actuators were continuously activated, the overall current would be $1460$ mA and the battery would last $5.62$ hours.
4.4. Maximum Number of Vibration Units On Each Chain
Here we described how the maximum number of units on each chain was determined (DR1). Due to incremental wire resistance along the chains, voltage drops will become more significant when additional vibration units are connected or when more units are activated simultaneously. To maintain stable operation, the MCU power line must sustain a minimum of $2.3$ V (i.e., the MCU’s minimum operational voltage) and the actuator power line should not fall below $0.9$ V (i.e., the LRA-X actuator’s rated voltage).
Figure 11. Voltage drops on the MCU power line and the actuator power line with an increasing number of actuators being activated.
[…]
This graph shows the voltage drop along the MCU and actuator power lines as the number of activated actuators increases, highlighting how the voltage decreases significantly for the actuators and slightly for the MCU, with critical thresholds marked for the MCU’s minimum operating voltage and the actuator’s rated voltage.
We measured the voltage drop threshold by incrementally activating vibration units on the chain. An oscilloscope was used to measure the MCU power line and actuator power line at the last vibration unit of the chain. If the last unit maintained a stable voltage, then one could be certain that other units were stable as well. Figure 11 shows the relationship between voltage drop number of actuators activated. When the number of actuators surpassed 16, the MCU voltage dropped below the operating range. When it surpassed 17, the actuator voltage dropped below its rated voltage. Therefore, control units could support up to 16 actuators on each chain.
4.5. Vibration Characteristics
To demonstrate that the toolkit can provide precise control over different vibration characteristics (DR2), we measured the acceleration of vibration units at various intensities and frequencies.
For intensity, we used an ADXL345 accelerometer for the measurements. Since most ADXL345 breakout boards are larger then the vibration unit itself, we designed a FlexPCB version and attached it between the skin and the vibration unit for precise measurements. A vibration unit was mounted on a sleeve and placed near the wrist of a human subject. During the testing, the vibration unit was activated for three seconds at each intensity level and the vibration frequency was kept constant at 170 Hz. The average acceleration was measured as the RMS value over the 1-second period of stable vibration. Acceleration was found to increase Monotonically as the intensity level increased (Figure 12 Left).
Figure 12. The acceleration measurements at various intensity and frequency levels.
[…]
This figure presents two measurements: on the left, a bar chart showing the increase in RMS acceleration as the intensity level rises from 1 to 15, and on the right, a frequency spectrum displaying the magnitude of accelerations at eight different frequency conditions ranging from 125 Hz to 384 Hz.
Similarly, for frequency, the vibration unit was placed near the wrist and activated for three seconds at each frequency level with level-3 intensity (25% duty cycle). To obtain the main frequency response, the acceleration data was processed using a fast Fourier transform. We found that the main frequency response was well aligned with the designated frequency at each level, thereby demonstrating the precise PWM signal control (Figure 12 Right).
4.6. Waveform Approximation
Because vibration units are controlled via discrete UART commands rather than continuous audio signals, one might worry that this would reduce the expressivity of the vibration waveform delivered through the units. Here we show that by utilizing high-bandwidth communication (DR2), VibraForge was capable of approximating audio signals and creating similar feedback. We used standard vibration waveforms from VibViz (Seifi et al., 2015), converted them from WAV files to HAPTIC files through Meta Haptics Studio (Meta Platforms, 2023a), and sent them to the GUI Editor. The editor automatically converted the haptic files to a sequence of vibration commands that were used to drive a vibration unit. We then measured the acceleration of the vibration unit and compared it with the original audio envelope. The results showed that they were closely approximated to each other (Figure 13).
Figure 13. Comparison of actuator vibrations (green lines) and original audio waveforms (light green shaded areas) from the VibViz Library
[…]
This figure compares actuator vibrations (green lines) and original audio waveforms (light green shaded areas) from the VibViz Library, showing four different waveforms.
5. Case Studies
We conducted three case studies to demonstrate the potential usage of VibraForge. The first case study showcased how VibraForge can be used to efficiently reproduce previous research prototypes. The second and third case studies showed how VibraForge can be used to build novel spatial vibrotactile systems that improve user performance in various applications.
5.1. Case Study 1: Replicating A Phonemic-Based Tactile Display
In Reed et al.’s work, a phonemic-based tactile display was designed to aid speech communication (Reed et al., 2018). This tactile display used 24 actuators on the forearm to deliver unique tactile codes of 39 English phonemes to deaf and hard-of-hearing users. Phonemes were mapped to vibration patterns using frequency, waveform, amplitude, spatial location, and movement. Although the display was shown to be useful, the process of creating the display was challenging. To drive the 24 actuators independently, the researchers had to use a desktop audio device (i.e., a MOTU 24Ao) with three custom amplifier boards to provide output signals, which made the whole system bulky and unable to be used in real-world situations. In this section, we demonstrated how the VibraForge toolkit could be utilized to simplify the development process and enhance the portability of the display.
5.1.1. Vibrotactile System Development
The first step in replicating the phonemic tactile display was to determine the layout of the actuators and the chain connections. Since the original display used a 4x6 layout on the forearm, we decided to use four chains with six vibration units on each chain and a small control unit.
Figure 14. The development process of Case Study 1’s phonemic tactile display re-implementation using VibraForge.
[…]
The top row shows all parts used for assembly, the assembly process itself, and the final prototype worn on an arm. The bottom row shows screenshots of waveform creation and layout design for the display, as well as an image of the original prototype used for comparison.
The second step was to assemble the display. The display required a sleeve garment and multiple modules from the toolkit (Figure 14). We first assembled the vibration units, then attached them to the garment and connected JST wires between them. The control unit was attached to the garment and four chains of vibration units were then connected to the control unit. The mapping between chains and connection ports followed the order of actuators described in the original paper. The whole assembly process took 16 minutes.
Once the hardware was built, the next step was to design the vibration patterns. Each phoneme needed to represented by a distinctive waveform that incorporated a high-frequency primary component and a low-frequency envelope, with a duration ranging from 100 to 480 ms. Consonants were designed as static patterns with multiple actuators being activated concurrently, whereas vowels were designed as dynamic patterns with actuators being activated in succession. The GUI Editor streamlined the process of creating the large variety of waveforms.
For example, to design the waveform for the consonant ”V” (IPA symbol /v/), we first dragged a sine wave into the Waveform Visualizer and set the frequency to be 300 Hz (primary component). Then we multiplied it with another sine wave at 8 Hz (envelope). The combined waveform was then saved in the Waveform Library as ”Consonant_B”. Alternatively, when the envelope was not available in the Waveform Library (e.g., $cos^{2}$ for consonant ”H”), we designed the waveform in Syntacts and imported the .CSV file into our GUI Editor. We then used the ”create new chain” function to create a virtual layout with four chains of vibration units on the canvas. We selected each of the designated vibration units (i.e., i6, ii6, iii6, iv6), assigned the custom waveform, and set the time range from 0 to 400 ms. Finally, we clicked the ”play” button to validate the design. The average time for an experienced haptic researcher on our research team to create one phoneme was one minute and five seconds. Therefore, the whole phoneme design process could be finished in 42 minutes. While the original paper did not mention the waveform design time, we believe the streamlined waveform creation process led to a significant reduction of design time.
5.1.2. Lessons Learned
The use of VibraForge to create the phonemic tactile display had several benefits. First, compared with the original device, our system was fully portable (DR5), with a small control unit and 24 vibration units directly attached to the sleeve. Second, VibraForge simplified the overall development process. The hardware system was reduced to module assembly without soldering, and the waveform design was supported by the GUI Editor (DR3). Third, the flexible design (DR4) enabled personalization opportunities. The original paper mentioned that the inter-motor spacing of the actuators needed to be adjusted for different participants and that the process of adjusting the layout was rather tedious. With our system, because the vibration units could be easily attached and detached, researchers could easily adjust the positions even if the garment was worn on the arm, thus giving them much more flexibility.
Meanwhile, some challenges were found when using the toolkit. While designers could easily author patterns using the GUI Editor, it mainly served as a testing site, because the current design of GUI Editor only allows users to manually click the ”play/stop” button to control the vibration sequences. This has constrained its use for real world interaction, where speech is segmented into phonemes and used to trigger corresponding patterns in real time. Right now the Python server does not support this kind of interaction so an additional API would be needed.
5.2. Case Study 2: Fitness Gaming in Virtual Reality
Virtual Reality (VR) fitness applications like OhShape (Lab, 2019) and Supernatural (Supernatural, 2024) have gained popularity by turning traditional fitness routines into interactive, gamified experiences. However, users often face challenges due to the absence of real-time, precise feedback about the physical positioning of their body (Wu et al., 2023). While prior research has demonstrated the utility of vibrotactile feedback for movement guidance in rehabilitation applications (Bark et al., 2014; Prabhu et al., 2020), systems used in these applications only deployed actuators on major joint locations such as the wrists or elbows, thus offering limited and coarse guidance. Inspired by OhShape (Lab, 2019) and the TV show ”Hole in the Wall” (Fox, 2008), we built a VR fitness application in which participants were asked to adjust their body poses to mimic cutouts in virtual walls when their digital avatar moved through the wall (Figure 15). We used the VibraForge toolkit to construct an upper-body suit that would provide spatial vibrotactile feedback to improve training performance.
Figure 15. An overview of the pipeline used in Case Study 2’s VR fitness game. Participant body movements were tracked by Vicon Motion Cameras and sent to a Unity scene for collision detection. Collision commands were then sent to a control unit on the participant’s back through Bluetooth and converted to UART commands for the vibration units.
[…]
The Vicon motion capture cameras captured one’s motion and transmitted the body tracking data to a Unity scene, where it detected if a collision occurred between the participant and a virtual wall. If a collision occurred, then it activated the corresponding actuator to deliver tactile feedback.
5.2.1. Vibrotactile System Development
The first step of our development was to determine the layout of vibration units on the body. We referenced prior work that measured spatial acuity (Huang et al., 2024a) and determined that vibration units should be uniformly distributed on the arm with 4 cm of spacing between them and on the body with 8 cm of spacing between them. As a result, the suit had 36 vibration units on the body and 40 vibration units on each arm (Figure 15).
The second step was to mount the control units and vibration units on the garment, which was an exercise compression shirt. Since the total number of vibration units was within the upper limit of the large control unit, we only needed to use one control unit. It was worn on the upper back to help users maintain their balance during fitness training. Vibration units were mounted one-by-one and JST wires were used to connect them.
The third step was to integrate the hardware into the fitness game, which was built using the Unity Game Engine (Technologies, 2020). The physical layout of the vibration units on the suit was manually converted into a virtual layout and overlaid on the user’s avatar, so that the virtual vibration units would move following the human body movement. Whenever a collision was detected between a virtual vibration unit and an obstacle, the system would determine which vibration units should be activated and send the vibration commands to the control unit through a Python server (Figure 15). Since collisions were binary events (i.e., collided / not collided), the vibration intensity was set to constant level (i.e., 7), as well as frequency (i.e., 170 Hz).
5.2.2. User Study Design
A preliminary study was conducted to evaluate the effectiveness of the vibrotactile suit on fitness training performance. The VR application was run on an Oculus Quest 2 headset (Meta Platforms, 2023b) through a Quest Link connection to a PC (i.e., Alienware M15 Laptop with RTX 3060Ti GPU). Vicon Motion Capture cameras (Vicon, 2023) were used to track the body poses of each participant and the poses were then mapped to the virtual avatar in the Unity scene. Nine participants (i.e., 4 Males and 5 Females; mean age = 25 years, SD = 3 years) were recruited to participate in the study, which lasted approximately one and a half hours, and each received \$40 compensation. The study was approved by our institutional research ethics board and conducted in a university lab space.
During the study, each participant would experience two conditions, i.e., no feedback and vibrotactile feedback. In the no feedback condition, participants only saw the incoming virtual walls and their own avatar and had to rely on proprioception to adjust their body position. In the vibrotactile feedback condition, participants would feel vibrations at the body positions where collisions occurred between vibration units and walls. For each condition, 24 pairs of walls with different cutouts were presented in sequence. Two walls in each pair were identical. The first wall served as the ”preview wall” to help the participant adjust their body pose, and the second wall served as the ”test wall” to measure any body position changes. We measured the number of collisions occurring during each condition. Our hypotheses were that (1) vibrotactile feedback would cause fewer collisions than no feedback, and (2) test walls would have fewer collisions than preview walls.
5.2.3. Results
The average number of collisions between vibration units and walls was computed for each feedback condition. Then, a repeated measures ANOVA was conducted using SPSS to determine if feedback condition had an effect on the number of collisions that occurred. If there were significant differences between each feedback condition, pair-wise comparisons with a Bonferroni correction were then performed.
The RM-ANOVA determined that both feedback type ($F(1,8)=6.450,p<0.05$) and wall type ($F(1,8)=5.582,p<0.05$) had significant effects on the number of collisions. No interaction effect was found (($F(1,8)=3.936,p=0.083$). For feedback type, the vibrotactile feedback (M=26.289, std=9.897) led to fewer collisions than the no feedback type (M=31.489, std=7.162, $p<0.05$). For wall type, the test walls (M=27.552, std=10) led to fewer collisions than the preview walls (M=30.226, std=7.694, $p<0.05$). Therefore, we concluded that both hypotheses were accepted.
5.2.4. Lessons Learned
The benefits of using VibraForge to build this high-density vibrotactile suit were multi-fold. First, scalability (DR1) enabled us to support a large number of actuators and provide highly localized feedback to guide body movement. Second, the modular and flexible design (DR4) made the system assembly process as simple as press-fitting the vibration units onto the garment and connecting wires between them. The entire process took less than 30 minutes. Compared with prior work that used sewing techniques (West et al., 2019; Lemmens et al., 2009), our solution was more flexible and efficient. Third, the low latency and high bandwidth communication (DR2) helped deliver almost real-time vibrotactile feedback to users when collisions occurred. Participants did not report any perceivable delays during the study.
Nonetheless, there were several challenges during the process. First, the integration of the vibrotactile system into the interactive game design was difficult. This was because we had to manually convert the physical layout of units on the body to a virtual layout in the Unity Engine and assign vibration units to their positions. This process would have been easier if we could detect the physical positions of the units and generate a virtual mapping automatically. Second, two participants reported that vibration units near the waist were not consistently perceived because body movement during fitness training would loosen the contact with garment. Our temporary solution was to tighten the suit so that movement was constrained. In the future, other methods would be needed to ensure that the vibration units maintained contact with the skin.
5.3. Case Study 3: Teleoperation Assistance for Unmanned Aerial Vehicle Operation
In this case study, we used VibraForge to design a wearable vibrotactile device to provide Unmanned Aerial Vehicle (UAV) operators with spatial information about obstacles. Current UAV teleoperation is dependent on visual feedback from onboard cameras, which presents significant challenges as a camera’s limited field of view can hinder situation awareness and lead to potential collisions with obstacles. While there have been prior attempts to use haptic feedback to notify operators of nearby obstacles (Brandt, 2009; Hou et al., 2017; Zhang et al., 2020; Xia et al., 2022), they required an operator to switch from using radio control (RC) controllers to hand-held haptic joysticks, only rendered force feedback in one direction at a time, or had fixed actuator layouts that could not represent lateral movements due to the absence of actuators on the side.
By creating a new vibrotactile feedback system using VibraForge, one could flexibly attach multiple vibration units to an operator’s body that would render simultaneous spatial cues about surrounding obstacles while the operator was using an RC controller. This would enhance operator’s collision avoidance ability and situational awareness in environments where visual feedback is limited.
5.3.1. Vibrotactile System Development
The first step of the development was to determine the layout of the vibration units on the body. Since there is no prior work that studied the mapping between tactile feedback positions on the body and human perception of 3D directions, we conducted a perceptual study with 10 participants to identify a mapping. The layout was initialized with 46 vibration units uniformly distributed on the body and optimized using user data. The final layout used 32 vibration units to achieve comprehensive coverage of the 3D space (Figure 16(b)). More details on the perceptual study and layout optimization can be found in (Huang et al., 2024b).
Figure 16. (a) The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point vibrotactile feedback on the body. Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange). (b) The layout of vibration units on the upper body. Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
[…]
(a) The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point vibrotactile feedback on the body. Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange). (b) The layout of vibration units on the upper body. Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
The second step was to affix control units and vibration units on the garment, which was a compression shirt similar to that used in Case Study 2. Thirty-two VCA vibration units were press-fit to the garment. A large control unit was placed on the operator’s back. The vibration units were grouped into four chains, each covering the front left, front right, back left, and back right.
The third step was to integrate the hardware into the simulated environment, which was built using Microsoft AirSim (Shah et al., 2017) with Unreal Engine 4 and Python APIs. Since the mapping between vibration units and 3D directions were known from the perceptual study, when a potential collision was detected, the system would trigger the unit that had the nearest direction with the colliding object in the space. This avoided the need to reproduce the physical layout of the vibration units in the system. The risks of colliding with surrounding obstacles were computed using control barrier functions (Ames et al., 2019) and scaled to $[0,15]$ to match the actuators’ intensity levels. If the risk of an obstacle was above $0.5$, the vibration unit in the corresponding direction would be activated by sending commands through a Python server.
5.3.2. User Study Design
A preliminary study was conducted to evaluate the effectiveness of the vibrotactile system on teleoperation performance. The simulation environment ran on an Alienware M15 laptop with RTX 3060Ti GPU. Twelve participants were recruited for the study (11 male, 1 female; mean age = 24 years, std = 3 years). Seven participants had previous experience operating quadrotors. Each study lasted 70 minutes and each participant received \$40 as compensation. The study was approved by our institutional research ethics board.
During the study, each participant went through three feedback conditions in a randomized order: no feedback (NA), force shared control (FSC), and vibrotactile shared control (VSC). For the NA and VSC conditions, participants used an Xbox controller for input, which had control mechanism similar to an RC controller. The output device for the VSC condition was the vibrotactile system. For the FSC condition, a Novint Falcon haptic joystick (Inc, 2024) was used for both input and output, similar to previous studies (Omari et al., 2013; Reyes et al., 2015).
For each condition, the experimental scene was a $5m\times 5m\times 50m$ tunnel (Figure 17(a)) that required participants to steer quadrotors forward, right, or upward. Different flying directions were used to vary visual information capacity as obstacles are often outside the camera’s field of view when flying right and upward. The tunnel contained fifteen objects that were randomly placed to avoid learning effects. Based on these conditions, we hypothesize that (a) vibrotactile feedback would enhance teleoperation performance compared to no feedback, and (b) the improvement would be more effective in the right and upward directions than the forward direction.
Figure 17. (a) A first-person view of the simulation environment used in Case Study 3’s study, with randomly positioned obstacles and a top-down view on the right. (b) The results for the total distance travelled, number of collisions, and input disagreement metrics for the feedback conditions (NA, FSC, VSC) and flying directions (forward-FWD, right-R and upward-UP). The error bars represent the standard error of the mean (SEM), $\mathchoice{{}^{\displaystyle*}}{{}^{\textstyle*}}{{}^{\scriptstyle*}}{{}^{\scriptscriptstyle*}}=p<0.05$, $\mathchoice{{}^{\displaystyle**}}{{}^{\textstyle**}}{{}^{\scriptstyle**}}{{}^{\scriptscriptstyle**}}=p<0.01$, and $\mathchoice{{}^{\displaystyle*{*}*}}{{}^{\textstyle*{*}*}}{{}^{\scriptstyle*{*}*}}{{}^{\scriptscriptstyle*{*}*}}=p<0.001$.
[…]
This figure shows an evaluation of a UAV teleoperation system. In (a), a navigation scene demonstrates the obstacle course where participants navigate using different control conditions. In (b), bar graphs display the performance metrics, including total distance, number of collisions, and input disagreement across multiple conditions (NA, FSC, VSC, etc.), highlighting significant differences between the control methods.
5.3.3. Analysis and Results
We evaluated three metrics: total distance travelled, number of collisions, and input disagreement (i.e., the deviation of user steering commands from the safe input range during flight). Using a two-way repeated measures ANOVA in SPSS, we assessed the effect of feedback condition and flying direction. Post-hoc pairwise comparisons with a Bonferroni correction were conducted for significant results (Figure 17(b)).
The RM-ANOVA determined that feedback condition had a significant effect on total distance travelled ($F(2,22)=3.585,p<0.05$), as did with flying direction ($F(2,22)=17.868,p<0.001$). The interaction between feedback condition and flying direction was also significant ($F(1.957,21.527)=4.916,p<0.01$). The post-hoc analysis of the interaction effect revealed that FSC-R had longer distance than FSC-FWD ($p<0.01$), FSC-UP ($p<0.01$), NA-R ($p<0.01$), and VSC-R ($p<0.05$), while VSC-FWD had a significantly shorter distance than VSC-R ($p<0.01$) and VSC-UP ($p<0.001$).
For the number of collisions, the RM-ANOVA determined that feedback condition had a significant effect ($F(2,22)=8.095,p<0.01$), as did flying direction ($F(2,22)=15.653,p<0.001$). The interaction between feedback condition and flying direction was also not found to be significant ($F(2.168,23.846)=1.910,p=0.168$). The pairwise comparisons between feedback conditions revealed that NA caused more collisions than FSC ($p<0.05$) and VSC ($p<0.05$). The pairwise comparisons of flying directions revealed that flying forward caused fewer collisions than right ($p<0.01$) and upward ($p<0.01$).
For input disagreement, the RM-ANOVA determined that feedback condition had a significant effect ($F(2,22)=4.798,p<0.05$), while flying direction did not ($F(1.254,13.789)=0.628,p=0.476$). The interaction between feedback condition and flying direction was also not significant ($F(4,44)=1.566,p=0.200$). The pairwise comparison between feedback conditions revealed that there was less disagreement with VSC compared to NA ($p<0.01$)
In summary, VSC caused fewer collisions and less input disagreement compared to NA, and achieved comparable performance with FSC. Additionally, the right and upward directions had more collisions than the forward condition, highlighting the increasing reliance users had on haptic feedback when under limited visual capacity. Thus, our hypotheses were accepted.
5.3.4. Lessons Learned
The process of using VibraForge to build the teleoperation vibrotactile system demonstrated several benefits. First, the flexible design (DR4) allowed us to iterate on the layout of vibration units rapidly during the perceptual study. From the initial 46 actuator layout to the final 32 actuator layout, we could easily adjust the positions of vibration units on the body and test various spatial mappings. Second, the low latency vibration responses (DR2) also provided instant feedback for participants, which is essential in high speed teleoperation tasks. In addition, the varied intensities (DR2) provided more nuanced information for participants, helping them prioritize their responses towards obstacles with higher risks. Together, these benefits demonstrate the usability of the toolkit for a complex application.
We also encountered several design challenges. One issue was that participants reported that the same intensity level was perceived differently across body areas. Locations with bone contacts were perceived to be stronger, while those with more subcutaneous tissues were perceived to be weaker. This variability may affect feedback precision during critical applications, and require a deeper investigation into how this can be overcome. Another issue was encountered during the layout iteration. Because the final layout was relatively sparse, the original connection wires that were used were too short. Therefore, we had to solder and assemble new wires. This, however, helped us make the toolkit adaptable to more diverse future use cases.
6. Discussion
In this section, we discussed the implications drawn from case studies, including strengths and limitations of VibraForge. Additionally, from our first-hand experiences in case studies, we synthesized a generalized five-step pipeline for building spatialized vibrotactile systems. Potential applications of VibraForge were highlighted at the end.
6.1. Strengths of VibraForge
VibraForge aims to provide a ”least resistance path” (Ledo et al., 2018) towards creating spatialized vibrotactile systems, which was demonstrated through the three case studies.
VibraForge offers great scalability. Systems designed in the case studies are of different scales, ranging from sleeve-type phonemic display with 24 vibration units on the forearm, to body-scale system with over 100 vibration units spanning across the upper body. Compared existing toolkits that are often constrained intrinsically to less than 16 actuators, VibraForge’s chain connection method enables designers to easily scale their system to more than 100 vibration units, significantly expanding the potential design space. Furthermore, adding a new unit to an existing system is also as simple as plug-and-play, without the need for any soldering or programming.
VibraForge offers great expressivity. Information of various complexity are adequately conveyed through spatial vibrotactile feedback in the case studies, ranging from simple on/off status for VR collision detection, to continuous risk levels in teleoperation and symbolic phoneme representation in the phonemic display. Compared with prior spatialized vibrotactile systems that are constrained to using ERMs and low-bandwidth control, our toolkit offers various actuator options and provides fine-grained control over vibration characteristics of each actuator. This enables vibration units to reliably reproduce audio waveforms in technical evaluations.
The GUI Editor of VibraForge features usability. In the phonemic display study, it effectively expedited the pattern authoring process by reducing the average pattern authoring time to one minute. Compared with prior design interfaces, our editor highlighted the canvas and the timeline that provided substantial support for multi-actuator pattern design.
VibraForge also features portability and flexibility. Compared with the original phonemic display that was wired to a bulky desktop audio device, components used in our prototype were more portable and easy to wear, such as the small control unit that communicates via Bluetooth. In the UAV teleoperation study, flexibility ensured that vibration units could be frequently adjusted during the layout optimization perceptual study. Compared to using commercial devices that often have fixed feedback positions, VibraForge enables designers to easily adjust the positions of vibrations units, thus enabling rapid design iterations.
6.2. Limitations of VibraForge
Although VibraForge can enable designers to quickly prototype and evaluate new multi-actuator vibrotactile devices and experiences, the toolkit does not come without limitations. For example, the open-loop control used in the chain connection topology does pose some issues. First, when a control unit sends a vibration command, it is unclear if the command was successfully transmitted to the target vibration unit or not. This is especially important when sending a ”stop” command, because the vibration unit won’t stop vibrating if the commands was missed during the transmission process. Second, when multiple actuators were activated, there was a noticeable voltage drop along the chains. Although the technical evaluation showed that the supply voltage remained in the MCU operating range and actuator voltage range, it would still affect the users’ subjective experience of the vibrations.
By adding extra command receiver pins on the control unit, one could instead use closed-loop control to solve these problems. With such a design, the vibration units and the control unit would form a loop, wherein the last vibration unit of the chain would be connected back to the control unit using a receiver pin. Whenever a vibration unit receives a command for itself, it would not only process the command and trigger the vibration, but also pass a confirmation message to the next unit. When the last unit in the chain receives this confirmation message, it would send the message back to the central unit, thus closing the loop. In cases where commands fail during the transmission process, the central unit would miss the confirmation message and resend the message. A closed loop connection would also reduce the voltage drop along the chain, as both ends of the chain would be connected to the power source. An LTspice simulation shows that an open-loop, eight-unit chain would result in voltage drop from 5 V to 3.1 V, while a closed-loop version would increase the voltage to 4.2 V.
Additionally, while the GUI Editor’s waveform design was advantageous for the phonemic display pattern design, designers would face difficulties when designing more complex spatial patterns. For example, if a designer wanted to create the effect of actuators vibrating from a central point and expanding out in concentric circles, a designer would have to manually design the waveform, assign the waveform to each vibration unit, and specify the start and end time, which is very tedious. A better solution would be to use graphical design metaphors and breakdown spatial patterns to low-level animations of point, line, and shapes. By leveraging phantom sensation illusions (Schneider et al., 2015), designers would only need to draw the shapes and trajectories, and they would be automatically converted to vibration commands for each unit. In this way, complex patterns could be created with ease.
Our case studies also emphasized the need for an automation pipeline to detect the positions of vibration units on the body and convert them to virtual layout that can be later used in various programming environment. We believe this need can be fulfilled with computer vision and motion capture techniques in the future. Using RGB images of the body, computer vision methods could detect the human body outline, segment vibration units from the body, and record the units’ positions. Similarly, retro-reflective markers could be mounted on top of each vibration unit to create a tracking object in the motion capture system. These positions could then be tracked in real-time and loaded into a development environment such as Unity Engine. In this way, designers could focus on the interaction design rather than manually creating the virtual layout.
6.3. Pipeline for Creating Spatialized Vibrotactile Systems
Synthesizing the processes followed during the three case studies, herein we summarize a generalized five-step pipeline for creating spatialized vibrotactile systems.
The first step is to determine what information to be delivered through vibrotactile feedback. For example, the purpose of the phonemic display was to convey vowels and consonants to deaf and hard-of-hearing people, so the types of vowels and consonants in words were the target information. In the drone teleoperation scenario, the key information was the potential risk of colliding with nearby obstacles and could be further divided into the directions of obstacles and their risk levels.
The second step is to decide how to convey the information. Since vibrations can be configured with different characteristics such as intensity, frequency, duration, position, and so on, the common practice is to breakdown the information and deliver it through multiple channels to maximize information transmission (Tan et al., 2020). For example, vowels and consonants for the phonemic displays were distinguished by using moving patterns and static patterns, while different consonants were distinguished by actuator position, frequency modulation and vibration duration. As shown in prior research, providing information through multiple channels with fewer levels in each channel can enhance the overall information transfer for vibrotactile systems.
The third step is to decide where to install the system and derive the layout of vibration units. In cases where the information was spatially connected to the surrounding environment, the layout should have a good mapping between tactile positions and spatial direction. For example, in the VR fitness application, vibration units were mainly placed on the arm to provide localized feedback, indicating the arm pose differences between the user and the wall. In the UAV teleoperation, locations of vibrations were perceptually mapped to the directions of incoming obstacle. Perceptual limits should also be considered, e.g., spatial acuity indicated the upper limits of layouts, as adding more actuators to the layout would not increase the information transmission.
The fourth step is to build the system. Control units and vibration units can be chosen based on the scale of the system. Learning from the building process of case studies, We recommend using the small control unit for localized systems such as sleeves and headbands, as well as applications that require short-term usage. Alternatively, the large control unit is suitable for long-term usage and body-scale systems, such as immersive virtual experience and robot teleoperation. When designing vibration patterns, if the information involves complex waveforms, designers can utilize the waveform library in the GUI Editor to expedite the authoring process, just like how we designed patterns for the phonemic display. If the information only requires changing levels of intensity or frequency, the Python server with APIs should be handy.
The final step is to test the system and iterate on the design. If the desired information cannot be transmitted as expected, the designer can reconfigure the system by changing the layout or experiment different characteristics for conveying the information.
6.4. Potential Applications for VibraForge
A toolkit such as VibraForge could benefit several domains. In VR, spatial vibrotactile feedback can enhance immersive social experiences. In VR social platforms, adding tactile interactions can help users express their emotions through social touch gestures (Sykownik and Masuch, 2020). If developers are creating devices for users with sensory issues in certain areas of their body, they can leverage the toolkit to quickly and easily adjust the layout of actuators to avoid discomfort.
During human-robot interaction, spatial feedback could be useful for assisting humanoid robot teleoperation. Humanoid robot training collects data from human operators performing tasks remotely. However, the absence of tactile feedback mechanisms makes it difficult for operators to execute daily tasks that are typically performed with ease using their own bodies (Pacchierotti and Prattichizzo, 2024). Spatial tactile feedback can provide additional information about remote interaction, such as when a robot hits a wall or other obstacle. The choice of tactile feedback positions can also be configured to match the positions of sensors on the robot. This feedback could thus increase an operator’s situational awareness and close the interaction loop.
Last but not least, we envision that VibraForge can be used to build custom prototypes for tactile perceptual studies, such as those measuring spatial acuity, information transfer, or other perceptual characteristics. For example, Park et al. performed perceptual studies to measure information transfer across five body positions (Park et al., 2023). These positions were not chosen by the researchers but instead were constrained by the design of the commercial system they used (i.e., tactile modules from bHaptics (bHaptics, 2024)). By using VibraForge, researchers could freely adjust the layout of vibration units on the body and adapt it to any test they want to conduct.
7. Conclusion
Acknowledging the limits with existing spatialized vibrotactile feedback systems and toolkits, we built the VibraForge toolkit. Our toolkit enables scalability via a modular design and the use of chain-connection data transmission methods. Technical evaluations validated the toolkit design, and case studies demonstrated the wide range of experiences that could be authored. The lessons learned during the case studies contributed to our understanding of a generalized pipeline for creating spatialized vibrotactile systems, and highlighted potential research directions of toolkit refinement and future applications. The toolkit is in its final stages of preparation, and will be released to the public later this year. We are excited to see how researchers and designers will harness the toolkit to elevate their own research and push the boundary of design innovation when using spatialized vibrotactile feedback.
Acknowledgements.

### v2

@@
 We also designed a GUI Editor to expedite the authoring of spatial vibrotactile patterns.
-Technical evaluations show that vibration units reliably reproduce audio waveforms with low-latency and high-bandwidth data communication.
-Case studies of phonemic tactile display, virtual reality fitness training, and drone teleoperation demonstrate the potential usage of VibraForge within different domains.
-Haptics, Vibrotactile, Toolkits, Wearable Devices
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Haptic devicesccs: Human-centered computing User interface toolkits
+Technical evaluation showed that vibration units reliably reproduced audio waveforms with low-latency and high-bandwidth data communication.
+Case studies of a phonemic tactile display, virtual reality fitness training, and drone teleoperation demonstrated the potential usage of VibraForge within different domains.
+A usability study with non-expert users highlighted the low technical barrier and customizability of the toolkit.
+Haptics, Vibrotactile Feedback, Haptic Toolkit, Multi-Actuator Devices, Wearable Devices
+journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714273isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Haptic devicesccs: Human-centered computing User interface toolkits
 Figure 1.
@@
 Leveraging a chain-connection method, one control unit in the system can independently control up to 128 vibration units.
-Applications of the toolkit are shown on the right.
+Applications of the toolkit include phonemic tactile displays, VR fitness training, and drone teleoperation.
 […]
-An overview of the VibraForge toolkit, that has control and vibration units, uses a chain-connection topology, and supports the creation of multi-actuator devices.
+An overview of the VibraForge toolkit, which has control units and vibration units, uses a chain-connection topology, and supports the creation of spatialized vibrotactile feedback systems.
 1. Introduction
-Tactile feedback conveys sensory information to the human body through the sense of touch (El Saddik et al., 2011).
-Vibrotactile feedback, which produces tactile sensations via actuators vibrating on the skin, has gained much popularity because it is compact, affordable, and accessible (Choi and Kuchenbecker, 2013).
-It has been widely adopted in various domains, such as virtual reality (VR) (Wang and et al., 2021; Kaul and Rohs, 2017; bHaptics, 2024), robotics (Pacchierotti and Prattichizzo, 2024; Xia et al., 2022; Aggravi et al., 2018), rehabilitation (Bark et al., 2014; Prabhu et al., 2020; Lindeman et al., 2006), and accessibility (Witteveen et al., 2015; Jung et al., 2022; Reed et al., 2018).
-To provide adequate feedback to users during complex tasks, spatialized vibrotactile feedback systems have been proposed (Lindeman et al., 2004; Louison et al., 2017).
+Tactile feedback conveys sensory information to the human body through the sense of touch (Weber, 1996).
+Vibrotactile feedback, which produces tactile sensations via actuators vibrating on the skin, has gained much popularity because it is compact, affordable, and accessible, as summarized by Choi et al.
+(Choi and Kuchenbecker, 2013).
+It has been widely adopted in various domains, such as virtual reality (VR) (Wang et al., 2021; Kaul and Rohs, 2017; bHaptics, 2024), robotics (Pacchierotti and Prattichizzo, 2024; SNAME, 2022; Aggravi et al., 2018), rehabilitation (Bark et al., 2014; Prabhu et al., 2020; Lindeman et al., 2006), and accessibility (Witteveen et al., 2015; Jung et al., 2022; Reed et al., 2018).
+To provide adequate feedback to users during complex tasks, spatialized vibrotactile feedback systems have been proposed, such as a battle simulation suit by Lindeman et al.
+(Lindeman et al., 2004) and directed movement sleeves by Louison et al.
+(Louison et al., 2017).
 These systems utilize actuator positions as an additional channel to enhance information transmission.
 For example, Wang et al.
-designed a facial mask with 36 eccentric rotating mass (ERM) actuators to study the perception of localization and orientation of objects in VR  (Wang and et al., 2021).
+designed a facial mask with 36 eccentric rotating mass (ERM) actuators to study the perception of localization and orientation of objects in VR  (Wang et al., 2021).
 The bHaptics TactSuit X40 (bHaptics, 2024) utilized 40 ERMs on a vest to create localized feedback for immersive VR gaming experiences.
 The bHaptics TactSuit was also used by Xia et al.
-to convey hydrodynamic flow intensity during submarine teleoperation (Xia et al., 2022).
+to convey hydrodynamic flow intensity during submarine teleoperation (SNAME, 2022).
 Jung et al.
@@
 Despite the great potential of spatialized vibrotactile feedback systems, previous implementations of such systems revealed scalability and expressivity limitations.
-Systems that depend on direct connections between actuators and microcontroller units (MCUs) exhibit limited scalability, as the number of actuators is intrinsically restricted by the available GPIOs or PWM channels on a MCU (Wang and et al., 2021; Kaul and Rohs, 2017; Reed et al., 2018).
+Systems that depend on direct connections between actuators and microcontroller units (MCUs) exhibit limited scalability, as the number of actuators is intrinsically restricted by the available GPIOs or PWM channels on a MCU (Wang et al., 2021; Kaul and Rohs, 2017; Reed et al., 2018).
 Expanding the number of actuators requires multiple MCUs, resulting in exponential increases in system complexity and communication overhead.
-For systems that use custom drivers and data bus protocols to drive actuators indirectly, the multi-layer connections support more actuators but compromise the vibration expressivity (LI et al., 2013; Lemmens et al., 2009; West et al., 2019).
+For systems that use custom drivers and data bus protocols to indirectly drive actuators, multi-layer connections support more actuators but compromise vibration expressivity (LI et al., 2013; Lemmens et al., 2009; West et al., 2019).
 These systems cannot render complex waveforms such as audio signals due to bandwidth limits, thereby limiting the types of actuators that can be used to simple actuators like ERMs.
 Moreover, while various toolkits have been developed to assist vibrotactile design, they typically employ direct or multi-layer connections and face the same limitations (e.g., TECHTILE (Minamizawa et al., 2012), VITAKI (Martínez et al., 2014), and VHP (Dementyev et al., 2021)).
-To overcome these limitations, this work proposes VibraForge, an open-source vibrotactile toolkit that leverages a scalable, modular design to enable the rapid prototyping of vibrotactile systems.
+To overcome these limitations, this work proposes VibraForge, an open-source vibrotactile toolkit [^1] The project resource can be found at: https://github.com/huangbj16/VibraForge that leverages a scalable, modular design to enable the rapid prototyping of vibrotactile systems.
 VibraForge consists of control units and vibration units.
-Control units are responsible for receiving vibration control commands from a Bluetooth server and sending them to vibration units.
+Control units receive vibration control commands from a Bluetooth server and sending them to vibration units.
 Vibration units are self-contained modules with an actuator driven by their own microcontroller.
-Vibration units are connected to a control unit via a chain-connection method, where a single GPIO pin on a control unit can control up to sixteen vibration units via a custom UART protocol.
-To better support spatial pattern authoring, we also designed an accompanying GUI Editor that enables haptic designers to intuitively author multi-actuator patterns.
-Technical evaluations of VibraForge showed that the toolkit has a high bandwidth (200 Hz), low latency (16 ms), and can reliably render complex audio signals in real time.
+They are connected to control units via a chain-connection method, where a single GPIO pin on a control unit can control up to sixteen vibration units via a custom UART protocol.
+To better support spatial pattern authoring, we also designed an accompanying GUI Editor that enables haptic designers to intuitively author multi-actuator tactile patterns.
+Technical evaluation of VibraForge showed that the toolkit has a high bandwidth (200 Hz), low latency (16 ms), and can reliably render complex audio signals in real time.
 Through three case studies, we demonstrate that VibraForge not only supports the efficient replication of existing research prototypes such as phonemic displays (Reed et al., 2018), but also improves user performance in various applications such as virtual reality fitness gaming and robot teleoperation.
-The contributions of this work are:
-  • an open-source hardware and software toolkit, VibraForge, that supports the design of spatialized vibrotactile feedback systems,
-  • technical evaluations of VibraForge, including power consumption, bandwidth, latency and acceleration, etc., and
-  • case studies demonstrating replicated research prototypes and novel applications enabled by VibraForge.
+In a usability study, ten participants with diverse technical backgrounds learned to use the toolkit in 15 minutes and completed design tasks and a free exploration activity.
+The findings from this study reaffirmed VibraForge’s low technical barrier and high potential for customizability.
+The contributions of this research are:
+  • an open-source hardware toolkit that supports the design of spatialized vibrotactile feedback systems,
+  • a novel chain-connection method that addresses the gap between system scalability and feedback expressivity,
+  • a software GUI editor that enables designers to intuitively author spatial vibration patterns,
+  • technical evaluations of VibraForge, including latency, bandwidth, vibration characteristics, etc., and
+  • case studies and a usability study demonstrating the usefulness and usability of the toolkit.
+Figure 2.
+Depiction of the direct connection (left) and multi-layer connection (middle) methods commonly used in spatialized vibrotactile systems and our proposed solution (right), which uses a chain connection method.
+[…]
+This figure illustrates three different connection methods for vibrotactile actuators: direct connection using PWM, multi-layer connection with SPI and custom drivers, and a chain connection solution using UART between vibration units for scalability.
 2. Related Work
@@
 Vibrotactile actuators such as eccentric rotation mass vibration motors (ERM), linear resonant actuators (LRA), voice coil actuators (VCA), or piezoelectric actuators (PIEZO) create mechanical vibrations via electromagnetic motions.
-Each actuator type and model has distinct voltage ratings, frequency responses, and vibration characteristics, making it challenging for designers to explore different configurations for wearable tactile devices.
-To assist the design process, various vibrotactile toolkits have been developed.
-These toolkits abstract low-level technical details and offer intuitive design interfaces, so researchers and designers can concentrate on creating vibration effects without being overwhelmed by the complexities of electronic, mechanical, and software design.
+Each actuator type and model has distinct voltage ratings, frequency responses, and vibration characteristics, making it challenging for designers to explore different configurations when creating wearable tactile devices.
+To assist their design process, various vibrotactile toolkits have been proposed.
+These toolkits abstract low-level technical details and offer intuitive design interfaces, so designers can concentrate on creating vibration effects without being overwhelmed by the complexities of electronic, mechanical, and software design.
 Vibrotactile prototyping toolkits can be categorized into those for education purposes and those for design purposes.
@@
 As vibrotactile feedback continues to be used for more complicated tasks and applications, these toolkits cannot meet the demands of emerging systems.
-Therefore, when designing VibraForge, we drew insights from previous spatialized vibrotactile feedback systems and propose a new connection method to overcome these scalability limitations.
+Therefore, when designing VibraForge, we drew insights from previous spatialized vibrotactile feedback systems and created a new connection method to overcome these scalability limitations.
 2.2. Spatialized Vibrotactile System Design
-Two connection methods have been often deployed When building spatialized vibrotactile systems (Figure 2).
-Some have tried maintaining a ”direct connection” to each actuator by adding more controllers to a system and thus increasing the number of PWM channels.
+Two connection methods are often used when building spatialized vibrotactile systems (Figure 2).
+Some have tried maintaining a ”direct connection” to each actuator by adding more microcontrollers to a system, thus increasing the number of PWM channels.
 For example, Wang et al.
-used PWM outputs from four Arduino Mega2560 boards to drive their facial display (Wang and et al., 2021).
+used PWM outputs from four Arduino Mega2560 boards to drive their facial display (Wang et al., 2021).
 Similarly, Kaul et al.
@@
 used three custom-built audio amplifier boards to send individual waveforms to 24 actuators on a sleeve to deliver speech phonemic cues to deaf users (Reed et al., 2018).
-Although complex waveforms could be sent to the actuators via direct connections, communication overhead and system complexity increased significantly when the system scaled up.
-Figure 2.
-Depiction of the direct connection (left) and multi-layer connection (middle) methods commonly used in spatialized vibrotactile systems and our proposed solution (right), which uses a chain connection method.
-[…]
-This figure illustrates three different connection methods for vibrotactile actuators: direct connection using PWM, multi-layer connection with SPI and custom drivers, and a chain connection solution using UART between vibration units for scalability.
+Although complex waveforms could be sent to the actuators via direct connections, communication overhead and system complexity increased significantly when the system was scaled up.
 Others have explored using ”multi-layer connections” that employ custom driver boards to interface between the microcontrollers and actuators (Sánchez and Bohne, 2015; LI et al., 2013; Lemmens et al., 2009; West et al., 2019).
@@
 Each PCB driver receives commands from the central MCU and generates PWM signals for its own actuator, thus supporting scalability.
-The high bandwidth and low latency control of vibration characteristics enables actuators to render complex waveforms, thus enhancing expressivity
+The high bandwidth and low latency control of vibration characteristics enables actuators to render complex waveforms, thus enhancing expressivity.
 2.3. Software for Vibration Pattern Design
@@
 For a comprehensive review of previous design software and their public availability, please see (Terenti and Vatavu, 2023).
-Here we discuss the common interfaces used to design single-actuator waveforms and multi-actuator patterns.
-For single-actuator design, one common method is to use predefined waveforms.
-Designers start with standard waveforms such as sine, square, or triangle, multiply them to generate new waveforms, and modify them using envelopes (Pezent et al., 2020).
+Here, we discuss the common interfaces used to design single-actuator waveforms and multi-actuator patterns.
+For single-actuator design, designers often start with standard waveforms (e.g., sine, square, etc), multiply them to generate new waveforms, and modify them using envelopes, as illustrated in Syntacts (Pezent et al., 2020).
 While this method provides authoring flexibility, it requires that designers have a deep understanding of how waveforms can be composed through combinations of standard signals.
@@
 VibraForge’s GUI Editor combines single-actuator waveform design with a multi-actuator canvas to support the efficient design for spatialized systems.
-It provides a library of standard waveforms and envelops for composing complex waveforms.
+It provides a library of standard waveforms and envelopes for composing complex waveforms.
 It is also compatible with existing tools such as Syntacts (Pezent et al., 2020) and VIREO (Terenti and Vatavu, 2023) to support the reuse of workflows and prior knowledge.
-The canvas shows a virtual layout of actuators that replicates the physical chain connections so designers can use it as reference when designing spatial patterns.
+The canvas shows a virtual layout of actuators that replicates the physical chain connections so designers can reference it while designing spatial patterns.
 3. VibraForge Design
-As the discussion of prior work highlighted, there is a trade-off between actuator expressivity and system scalability.
-We aimed to mitigate this trade-off by using a novel chain-connection method with custom designed vibration units and control units.
-This section provides an overview of the VibraForge toolkit [^1] The project resource can be found at: https://anonymous.4open.science/r/VibraForge, including its design requirements, vibration units, control units, chain-connection method, signal segmentation algorithm, and vibration pattern GUI Editor.
+As highlighted in prior work, a trade-off exists between actuator expressivity and system scalability.
+To address this, we developed a novel chain-connection method with custom vibration units and control units.
+This section provides an overview of the VibraForge toolkit, covering its design requirements, components, chain-connection method, signal segmentation algorithm, and GUI Editor.
 3.1. Design Requirements
-VibraForge’s design requirements were drawn from design challenges discussed in prior work on wearable spatial vibrotactile design or were identified by the authors based on prior prototyping experiences.
-From spatial vibrotactile design, three design requirements were proposed:
-  • DR1: Scalability. The toolkit should support large quantities of actuators. Designers should be able to use different types of actuators such as ERM, LRA, and VCA.
+VibraForge’s design requirements were drawn from design challenges discussed in prior work on spatial vibrotactile and wearable design, or identified by the authors based on prior prototyping experiences.
+  • DR1: Scalability. The toolkit should support large quantities of actuators such as ERM, LRA, and VCA.
   • DR2: Fine-Grained Control.
@@
   • DR3: Usability.
-The toolkit should provide an intuitive design interface for authoring vibrotactile spatial patterns.
-To leverage designers’ prior design knowledge, the toolkit should be compatible with existing vibrotactile waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a).
-From wearable design, two design requirements were proposed:
+The toolkit should provide an intuitive design interface for authoring vibrotactile spatial patterns that leverages designers’ prior design knowledge and is compatible with existing vibrotactile waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a).
   • DR4: Flexibility.
-The toolkit should enable designers to modify settings, reprogram behaviors, and physically rearrange components without difficulty.
-This requirement was derived from design challenges faced by West et al.
-when attempting to relocate or replace components that were sewn onto a garment (West et al., 2019).
+The toolkit should enable designers to modify settings, reprogram behaviors, and physically rearrange components without difficulty, derived from design challenges faced by West et al.
+(West et al., 2019).
   • DR5: Portability.
-Devices built with the toolkit should be portable and powered by batteries, communicate with external devices via wireless protocols, and use lightweight components to ensure a comfortable wearing experience.
+Devices built with the toolkit should be powered by batteries, communicate with external devices via wireless protocols, and use lightweight components to ensure a comfortable wearing experience.
 3.2. Vibration Units
-The vibrations units were designed as self-contained modules that could drive actuators by themselves.
-Each vibration unit consisted of a vibrotactile actuator, a custom designed PCB, and 3D printed components (Figure 3).
-We designed three versions of the vibration unit to support a range of actuators (DR1).
+The vibrations units were self-contained modules comprised of a vibrotactile actuator, a custom designed PCB, and 3D printed components (Figure 3).
+Three versions were designed to support various actuators (DR1).
 Figure 3.
 (Left) Exploded views of the small vibration unit for X-axis LRAs, including the cap, the PCB, the PCB mount, the enclosure, and the bottom ring.
-Z-axis LRAs shared most components except the enclosure.
+Z-axis LRAs shared most components except for the enclosure.
 (Right) Exploded view of the large vibration unit for VCAs with a similar structure.
@@
 3.2.1. Vibrotactile Actuators
-An X-axis, square-shaped, 9.5 mm x 9.5 mm LRA manufactured by the Jinlong Machinery and Electronics Company (Model L959535) (Machinery and Co., 2022) was chosen for the vibration unit.
-Its resonant frequency was 170 Hz.
-When powered by 0.9 Vrms, it had a maximum current of 145 mArms and an acceleration of 0.65±0.12 Grms.
-The other two actuators were a Z-axis LRA manufactured by Jinlong (Model G0832) (Machinery and Co., 2022) and a VCA manufactured by PUI Audio, Inc.
-(Model HD-VA3222) (PUI Audio, 2023).
-The Z-axis LRA was cylindrical-shaped, with an 8 mm diameter and a height of 3.2 mm.
-Its resonant frequency was 235 Hz.
-When powered by 1.8 Vrms, it had a maximum current of 90 mArms and an acceleration of 0.50±0.08 Grms.
-The VCA was 32 mm x 22 mm and had a frequency response range from 80 - 500 Hz.
-When driven by a 133 Hz 1.5 Vrms signal, it had a maximum acceleration of 2.52 Gp-p.
-These actuators were chosen because they cover a variety of vibration directions and intensities, thus satisfying the tactile perception requirements of different body parts (i.e., spatial acuity (Elsayed et al., 2020) and perceived intensity (Verrillo et al., 1969)).
-Since they all share the same PCB schematics and similar enclosure structures, in the following sections we only describe the vibration unit designed for the X-axis LRA.
-The other designs can be seen in Figure 3.
+The vibration unit used a square-shaped X-axis LRA (9.5 mm x 9.5 mm, Model L959535, Jinlong Machinery and Electronics (Machinery and Co., 2022)), resonating at 170 Hz with a maximum current of 145 mArms and max acceleration of 0.65±0.12 Grms at 0.9 Vrms.
+Two additional actuators included a cylindrical Z-axis LRA (8 mm diameter, 3.2 mm height, Model G0832, Jinlong (Machinery and Co., 2022)) resonating at 235 Hz with 0.50±0.08 Grms of acceleration and a maximum current of 90 mArms at 1.8 Vrms, and a VCA (32 mm x 22 mm, Model HD-VA3222, PUI Audio Inc.
+(PUI Audio, 2023)) with a 80 - 500 Hz response range and 2.52 Gp-p acceleration at 133 Hz, 1.5 Vrms.
+These actuators covered various vibration directions and intensities to meet the tactile perception needs of different body parts (i.e., spatial acuity (Elsayed et al., 2020) and perceived intensity (Verrillo et al., 1969)).
+Since they share the same PCB schematics and similar enclosures, we will focus on the X-axis LRA vibration unit in the following sections, with the other units shown in Figure 3.
 3.2.2. Printed Circuit Board Design
-Due to bandwidth limitations, as the number of actuators increases it becomes difficult to drive them using real-time PWM signals sent from a central unit.
-Therefore, we allocated an MCU to each vibration unit to drive each actuator.
-Each vibration unit had its own driver PCB to receive UART commands from a central board and provide fine-grained control over vibration characteristics.
-The driver PCB had a PIC16F18313 MCU, a DRV8837 H-bridge motor driver, two 4-pin JST-SH headers, two pogo pins, and supportive components (Figure 4).
-The MCU was used to transmit commands and generate waveforms.
-The JST headers were used as input and output for the power and signal connections.
-The number of connector pins was minimized to four (i.e., actuator power, MCU power, ground, and data).
-The pogo pins were mounted on the bottom to ensure direct contact with the actuator pads.
-Figure 4. The circuit schematics and PCB board designs of vibration units.
+Each vibration unit had a PCB driver board to receive UART commands from a control unit and provide fine-grained control over vibration characteristics (DR2).
+The PCB had a PIC16F18313 MCU, a DRV8837 H-bridge driver, two 4-pin JST-SH headers, and two pogo pins for actuator connections (Figure 4).
+Upon receiving a UART command from the RX line of the JST connector, the MCU extracted vibration characteristics from the command, generated waveform using an internal PWM generator, and sent it to the actuator through the H-bridge driver.
+The communication protocol is detailed in Section 3.4.
+Figure 4. The circuit schematics and PCB board designs of the vibration units.
 […]
 Circuit schematics of a unit driver that generated two opposite waveforms based on the received control signals to modulate the actuator.
-When the MCU received a ”start” signal, it sent a waveform to the H-bridge to modulate the actuator.
-We utilized the PWM generator inside the MCU for fine-grained control over the actuators’ intensities, frequencies, and waveforms.
-Because there is no address system for each unit, designers only need to load their program into the MCU once.
-They can then plug-and-play each unit as desired.
-The communication protocol is further described in Section 3.4.
 3.2.3. 3D Printed Enclosures
-The 3D printed enclosures were designed to enclose each vibrotactile actuator and PCB board and provide a method to easily attach and detach each vibration unit to a garment (DR4).
-Similar to the size of a coin, the enclosure was compact and lightweight, enabling for a comfortable wearing experience (DR5).
-The enclosure consisted of four parts: a cap, a PCB mount, an enclosure, and a bottom ring (Figure 3).
-The cap employed a screw-nut fastening system, which added top-down pressure to stabilize the PCB.
-The PCB mount offered protection for the PCB and aligned the pogo pins directly with the actuator pads.
-This improved the robustness of the vibration unit as no soldering was required during assembly.
-The enclosure served as a protective housing for all the parts.
-It had two holes for JST connectors, as well as two asymmetrically positioned vents for passive heat dissipation.
-The bottom ring was designed to be positioned beneath the garment to maintain direct contact with the skin without the need for any sewing (DR4).
-The same design can be easily adapted to other vibration unit structures for other types of actuators.
-In Figure 3, other vibration units that share the same PCB schematics and slightly different enclosure designs are shown.
+The 3D printed enclosures were designed to enclose actuators and PCBs and allow easy attachment to garments (DR4).
+Compact and lightweight like a coin, they ensured a comfortable wearing experience (DR5).
+The enclosure included a cap with a screw-nut system to stabilize the PCB, a mounting plate for PCB protection and alignment of the pogo pins, a protective housing with two JST connector holes and heat dissipation vents, and a bottom ring that maintained direct skin contact without sewing (DR4) (Figure 3).
+The same design strategy was also applied to design enclosures for other actuators.
 3.3. Control Unit
-The control unit was designed to be fully portable (DR6).
-It consisted of a PCB extension board with an ESP32 microcontroller (MCU), two LiPo batteries, and a 3D-printed enclosure (Figure 5).
-Multiple versions of control units were designed to accommodate different projects, ranging from simple arm bands to more spatialized body displays.
-They varied in the (a) number of chains supported, (b) model of the ESP32 board, and (c) battery capacity.
+The portable control unit (DR5) featured a PCB extension board with ESP32 MCU, two LiPo batteries, and a 3D-printed enclosure (Figure 5).
+Multiple versions were designed to accommodate different project requirements, varying in the number of supported chains, the ESP32 model, and battery capacity.
 Figure 5. The control unit, which contained an ESP32 MCU, an extension board, chain connectors, and LiPo batteries.
@@
 This figure shows an ESP32 board mounted on the extension board in the middle of the control unit, and the 8 chain connection ports extending out from the control unit.
-We designed a large control unit that supported up to 8 chains of vibration units and 16 units per chain, for a total of 128 vibration units.The extension board drew power from two serial 3.7 V 8200 mAh LiPo batteries.
-It stabilized the voltage at 5 V through two power regulators: a linear regulator that powered all the MCUs and a buck regulator that powered all actuators in the system.
-These separate power supplies ensured that MCUs operated on stable voltages without interference from the actuators’ power consumption.
-The extension board had eight JST connectors.
-Each connector was connected to a GPIO pin on the ESP32 MCU and drove a chain of vibration units.
-The MCU was an Adafruit Feather ESP32 V2222https://learn.adafruit.com/adafruit-esp32-feather-v2.
-It was responsible for converting external Bluetooth commands to UART commands and sending them to the corresponding vibration unit.
-Since the number of required vibration units varies across different tasks, we also designed a small control unit with fewer supported chains.
-This compact unit design contained two 3.7 V 500 mAh LiPo batteries and an Adafruit QT Py ESP32-S3  [^3] https://learn.adafruit.com/adafruit-qt-py-esp32-s3/pinouts.
-This MCU had four general purpose GPIO pins that could be used for chain connections.
+We designed a large control unit supporting up to 8 chains, with 16 units per chain (128 total).
+Powered by two 3.7 V 8200 mAh LiPo batteries, the unit used a linear regulator for the MCUs and a buck regulator for the actuators, ensuring 5 V stable voltage supplies for both.
+The extension board had eight JST connectors for vibration units, each linked to a GPIO pin on an Adafruit ESP32 Feather V2 MCU  [^2] https://learn.adafruit.com/adafruit-esp32-feather-v2.
+The MCU converted Bluetooth commands to UART commands and sent to corresponding vibration units.
+For devices requiring fewer units, we also created a more compact control unit with two 3.7 V 500 mAh LiPo batteries and an Adafruit QT Py ESP32-S3 MCU  [^3] https://learn.adafruit.com/adafruit-qt-py-esp32-s3, supporting four chain connections.
 3.4. Chain-Connection Method and Data Communication
-VibraForge differs from existing toolkits due to its use of a chain-connection method (Figure 6).
-This method enables the toolkit to address over 100 vibration units using a single control unit (DR1) and support the simultaneous control over the vibration characteristics of multiple actuators (DR2).
+VibraForge differs from existing toolkits by using a chain-connection method (Figure 6), allowing a single control unit to manage over 100 vibration units (DR1) and simultaneously control multiple actuators’ vibration characteristics (DR2).
 Figure 6. The multiple chain-connection method used within VibraForge, depicting the connections required for a given chain.
@@
 This figure shows where the control unit, the power converters, and unit drivers are located to illustrate the multiple chain-connected method.
-As explained in Section 3.3, each GPIO pin on the ESP32 MCU was routed to a JST connector on the extension board.
-Because vibration units shared the same JST connectors, they could be connected in chains and receive commands through the same GPIO pin.
-There were two options for the data transmission design.
-The first option was to use data bus protocols such as CAN or I2C, where all the vibration units would receive the same command simultaneously and use the command’s address to determine whether the command was intended for itself.
-While this option would allow commands to be transmitted with a minimal latency, the vibration units would need to be programmed with different addresses prior to use, which would be burdensome when working with multiple units.
-Instead, we decided to send commands through each chain using a serial protocol, where each unit would receive a command, process it, and resend it to the next unit.
-While this may increase the overall latency, it allowed each unit to operate with a ”virtual address”, enabling independent plug-and-play control without additional programming.
-Figure 7. The two-byte message bits definition and an example of how vibration commands are transmitted along a chain using the ”virtual address” system.
+Vibration units are connected in chains to receive commands from the ESP32 GPIO pins.
+We considered two data transmission options (a) data bus protocols (e.g., CAN or I2C) would guarantee minimum latency via broadcasting commands, but require programming different addresses for each unit; (b) serial protocols (e.g., UART) did not require unique address, but passing commands along the chain would increase latency.
+We decided to use UART because it enabled plug-and-play control without extra programming (DR4) and latency could be reduced to imperceptible levels via high-speed transmission.
+Vibration commands were transmitted via UART at 115.2 kHz with parity checks (Figure 7).
+Each command included two bytes: the first byte encoded the address and start/stop status, and the second byte specified vibration characteristics.
+Upon receiving the first byte, a unit checked the address, decremented it, and forwarded the command to the next unit.
+This repeated until the address was zero, thus signaling the target unit had been reached.
+The targe unit acted based on the start/stop bit: ’0’ (STOP) disabled the actuator, while ’1’ (START) activated it.
+The second byte detailed the vibration intensity (16 levels from 0% to 100% ) and frequency (i.e., 8 levels following Weber’s factor for frequency differences (Pongrac, 2006), i.e., 123, 145, 170, 200, 235, 275, 322, and 384 Hz).
+Figure 7. The two-byte message bits definition and an example of how vibration commands were transmitted along a chain using the ”virtual address” system.
 […]
@@
 The address and start/stop bit were in the first message, and operational commands including the duty cycle, frequency, and waveform, were in the second message.
-Vibration commands were transmitted via a UART protocol at 115.2 kHz with parity checks (Figure 7).
-Each command consisted of two bytes: the first byte encoded the address and start/stop status of the target vibration unit, and the second byte detailed the vibration characteristics.
-Upon receiving the first byte, a vibration unit checked the address.
-A non-zero address indicated the command was for subsequent units, prompting the unit to decrement the address and forward the message.
-This process continued until the address reached zero and signified the targeted unit, which then acted based on the start/stop bit: ’1’ (START) awaited the second byte, while ’0’ (STOP) disabled its actuator.
-When a unit was ”signified”, it waited for the second-byte message, which contained details about vibration characteristics, including the 16 intensity levels (evenly distributed from 0% to 100% duty cycle) and 8 frequencies (i.e., 123, 145, 170, 200, 235, 275, 322, and 384 Hz).
-The frequency levels were chosen so that each level was 1.18 times higher than the previous level.
-This followed Weber’s factor for vibrotactile frequency just-noticeable differences (Pongrac, 2006).
 3.5. Signal Segmentation Algorithm
-To render complex waveforms such as audio signals on vibration units (DR2), we implemented a signal segmentation algorithm that converted high-frequency input signals to low-frequency serial commands (Figure 8).For instance, consider an input waveform of 44100 Hz audio signals that composed of 200 Hz and 5 Hz sine waves.
-High-frequency components (200 Hz) in the waveform were extracted using a short-time Fourier transform (STFT), which captured how the dominant frequency changed over time.
-Low-frequency components (5 Hz) were extracted using a Hilbert transform, which captured the envelope of the waveform.
-Both components were downsampled to 200 Hz and sent to vibration units.
-Since our frequency levels ranged from 123 Hz to 384 Hz, the segmentation threshold was set to 100 Hz.
-Frequency components above 100 Hz were rendered using frequency levels, while components below 100 Hz were rendered using intensity levels.
-The technical evaluations in Section 4 demonstrated that this algorithm could reliably render complex waveforms such as audio signals using serial command control.
+To render complex waveforms like audio signals (DR2), we implemented a signal segmentation algorithm to convert high-frequency input signals to low-frequency serial commands (Figure 8).
+For example, a 44100 Hz audio signal with 200 Hz and 5 Hz sine waves would be processed using a short-time Fourier transform (STFT) to extract high-frequency components (main frequency) and a Hilbert transform for low-frequency components (envelope).
+Both components were then downsampled to 200 Hz and sent to vibration units.
+Components above 100 Hz would be transmitted via frequencies, and those below 100 Hz as intensities.
+Technical evaluation (Section 4) confirmed the algorithm ’s reliability in rendering complex waveforms.
 Figure 8. The signal segmentation workflow, showing how low frequency and high frequency components were extracted from a complex waveform.
@@
 3.6. GUI Editor for Designing Spatialized Patterns
-We designed a GUI Editor (Figure 9) to enable designers to create spatialized vibration patterns (DR3).
-This editor differed from previous editors in that it combined the waveform editing from single-actuator design software with the canvas design used for multi-actuator design.
-The key components of the GUI Editor were:
+We designed a GUI Editor (Figure 9) to support the creation of spatialized vibration patterns (DR3).
+The editor combined the waveform editing from single-actuator design software with the canvas design used for multi-actuator design.
+Its key components were:
 Figure 9. The layout of the GUI Editor, with its Waveform Library, Waveform Visualizer, Canvas, and Timeline.
@@
 The Timeline at the bottom displayed the sequence and timing of waveform actions and events.
-  • Waveform Library: The Waveform Library contained waveform templates such as oscillators and envelopes.
-Designers could use them to create a custom waveform and save it to the Waveform Library.
-In addition, designers could import waveforms in the JSON format from popular waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a) so they could reuse previous designs (DR3).
-  • Waveform Visualizer: The Waveform Visualizer served two purposes.
-When designers were creating a custom waveform, they could drag and drop waveform templates from the Waveform Library onto the Waveform Visualizer and it would show the combined waveform.
-When designers were composing vibration patterns for individual vibration units, the Waveform Visualizer also show the composed waveform for the selected unit.
-  • Canvas: The Canvas enabled designers to drag-and-drop vibration units onto the canvas so they could be connected into chains that mimicked the physical layout of the vibration units.
-In addition, a ”Create New Chain” function enabled designers to generate vibration units in batches and arrange them using grid layouts.
-When clicking on a unit, its corresponding waveform would show up in the Waveform Visualizer.
-  • Timeline: The Timeline served as a high-level summary of the vibration patterns.
-The blue blocks showed the range of time when the vibration unit was activated.
-By moving the slider left and right, the vibration units that were activated at that moment would be highlighted in the Canvas so designers could understand the vibration patterns over time.
-By clicking the ”Play” button, the waveform would be converted into vibration commands with timestamps and sent to the control unit via Bluetooth.
-In a typical spatial pattern design workflow, designers would begin by creating a new design and specify the number of chains and actuators to be used in their system.
-The GUI would then generate a default Canvas with the designated number of actuators.
-Designers would then freely adjust the layout to match their physical device.
-Then, designers would assign waveforms to each vibration unit and specify the start and end time.
-Each waveform could be designed from scratch using the waveform templates in the Waveform Library, or imported from other software.
-To review the spatialized patterns, designers would click on any vibration unit on the Canvas and the Waveform Visualizer would show the corresponding waveform.
-Alternatively, designers could see a high-level summary of how the vibration patterns of multiple units would change over time by moving the Timeline slider.
-Additionally, we also provided a Python server with VibraForge APIs, so that vibration units can be directly activated and controlled via code (DR3).
-Later in Section 5, Python servers are integrated into external programming environments (e.g., Unity) to trigger real-time feedback for applications.
+  • Waveform Library: The Waveform Library included templates like oscillators and envelopes for creating custom waveforms.
+Designers could also import JSON waveforms from editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a), enabling for the reuse of previous designs (DR3).
+  • Waveform Visualizer: The Waveform Visualizer allowed designers to create custom waveforms by combining templates from the Waveform Library via drag-and-drop.
+It also displayed the composed waveform for selected vibration units during pattern composition.
+  • Canvas: The Canvas provides an approximate virtual layout of vibration units and their relative order within each chain.
+Designers could drag-and-drop a single unit into the Canvas.
+A ”Create New Chain” function also allowed them to create a new chain of units at once and align the units in grid patterns.
+Selecting a unit would display its waveform in the Waveform Visualizer.
+  • Timeline: The Timeline summarized vibration patterns, with blue blocks indicating activation periods.
+Moving the slider highlighted active units on the Canvas.
+Clicking the ”Play” button sent real-time vibration commands to the control unit via Bluetooth.
+In a typical spatial pattern design workflow, designers would specify the number of chains and actuators for their system.
+The GUI then generated a default Canvas with the designated number of actuators, which could be freely adjusted to match their physical layout.
+Designers then assigned waveforms to each vibration unit, specifying start and end time.
+Waveforms can be created using templates from the Waveform Library or imported from other software.
+To review patterns, designers could click on a vibration unit to view its waveform in the Waveform Visualizer or use the Timeline slider to see a high-level summary of vibration patterns over time.
+This workflow was found to be easy to follow by both experts (Section 5) and novice users (Section 6).
+To support the integration of VibraForge with external programming environments (DR3), we created a Python server and a Unity plugin that provided APIs to directly address the VibraForge hardware.
+The APIs offered functions such as SendCommand(address, start_stop, intensity, frequency) that enabled designers to configure the vibration parameters of a unit without delving into the waveform details, thereby assisting users without extensive technical expertise.
+The APIs were later used in Case Study 2 and 3 to build VR and robotics applications (Section 5).
 4. Technical Evaluation
-In this section, we summarize various performance metrics of the VibraForge toolkit to demonstrate how it fulfills the aforementioned design requirements.
-All tests below used vibration units of X-axis LRAs and the small control unit with default level-7 intensity (50% duty cycle) and 170 Hz frequency, unless otherwise stated.
+In this section, we demonstrate how the VibraForge toolkit fulfills the aforementioned design requirements via various performance metrics.
+For the evaluations, we used vibration units of X-axis LRAs and the small control unit with a 50% intensity and frequency of 170 Hz, unless otherwise noted.
 4.1. Latency
-Having fine-grained control (DR2) over the actuators in a vibrotactile system means that the data transmission of the system must have low latency and high bandwidth.
-To determine if the latency of VibraForge was below the human perceivable latency threshold of 45 ms (Vogels, 2004), we measured the end-to-end latency of a PC to a control unit (i.e., the transmission latency of the Bluetooth Low Energy protocol (BLE)), and of a control unit to a vibration unit (i.e., the latency of the custom UART protocol).
-To measure the PC to control unit latency, a BLE write GATT command was sent and immediately followed by a BLE read GATT command.
-A Python timer recorded the write command sending time and read command receiving time and computed the difference as the latency.
-We repeated this measurement 100 times and found that the average round-trip transmission latency was $26.67\pm 6.53$ ms.
+Low latency and high bandwidth are crucial for data transmission in spatialized vibrotactile systems, particularly in scenarios where user actions trigger immediate tactile feedback, enabling users to respond and adjust their behaviors accordingly.
+To determine if VibraForge’s latency was below human perceivable threshold of 45 ms (Vogels, 2004), we measured the end-to-end latency of a PC to a control unit (i.e., latency of the Bluetooth Low Energy protocol (BLE)), and of a control unit to a vibration unit (i.e., latency of the UART).
+To measure the PC to control unit latency, a BLE GATT write command was sent and immediately followed by a read command.
+A Python timer computed the difference between the write command sending time and read command receiving time.
+We repeated this measurement 100 times and found the average round-trip transmission latency was $26.67\pm 6.53$ ms.
 Therefore, the one-way transmission latency was approximately $14$ ms.
-For the control unit to vibration unit latency, repeated measurements using an oscilloscope showed that each command took $125$ us to be transmitted from one unit to the next unit on a chain.
-Therefore, latency on the longest chain with 16 actuators would accumulate to $2$ ms.
-Adding the two measurements together would thus result in an overall latency of approximately $16$ ms, which would be virtually imperceptible by humans (Vogels, 2004).
+For the control unit to vibration unit latency, repeated measurements using an oscilloscope showed that each command took $125$ us to be transmitted from one unit to the next unit.
+Therefore, latency on a chain with 16 actuators would be $2$ ms.
+These added to an overall latency of approximately $16$ ms, which would be virtually imperceptible by humans (Vogels, 2004).
 4.2. Bandwidth
-The bandwidth of VibraForge was the number of commands that could be sent and processed per second by control units and vibration units.
+VibraForge’s bandwidth was defined as the number of commands that could be sent and processed per second by the control and vibration units.
 Bandwidth differs from latency in that asynchronous BLE commands could be continually sent without needing to wait for the previous command to be received (Figure 10).
-Therefore, bandwidth represented the refresh rate of the vibration commands for each actuator.
-Since BLE commands (millisecond level) would be processed much slower than UART commands (microsecond level), the bandwidth bottleneck of the system depended on the processing speed of BLE commands on the control unit.
+Since BLE commands (i.e., millisecond level) would be processed much slower than UART commands (i.e., microsecond level), the bandwidth bottleneck depended on the processing speed of the BLE commands on the control unit.
 Figure 10. Data transmission bandwidth breakdown.
 […]
-This figure shows the breakdown of data transmission bandwidth and latency, highlighting BLE (Bluetooth Low Energy) transmission and UART communication.
-The BLE transmission segments are separated by 14 milliseconds, followed by 2 milliseconds of UART communication.
-The bandwidth is marked as 5 milliseconds or 200 Hz, representing the frequency of data transmission cycles.
-The diagram visually represents how BLE and UART transmissions alternate in time.
-To measure VibraForge’s bandwidth, we wrote a Python testing program to repetitively send standard BLE commands without latency, and another program on the ESP32 (i.e., control unit) to receive the commands and print the receiving time to the serial port.
-The command processing time was computed as the difference between the receiving time for each command.
-We sent $1,000$ BLE commands with five UART commands in each of them, and computed an average processing time as $2.96\pm 1.20$ ms.
-Therefore, for the actual operation of the toolkit, we set the BLE command sending delay to be $5$ ms with some tolerance.
-This means that the control unit could simultaneously update five vibration units at $200$ Hz refresh rate, satisfying the signal segmentation algorithm requirements.
+The breakdown of data transmission bandwidth and latency, highlighting BLE (Bluetooth Low Energy) transmission and UART communication.
+The BLE transmission segments were separated by 14 milliseconds, followed by 2 milliseconds of UART communication.
+The bandwidth was 5 milliseconds or 200 Hz, representing the frequency of the data transmission cycles.
+The diagram represents how BLE and UART transmissions alternate in time.
+To measure VibraForge’s bandwidth, a Python testing program repetitively sent BLE commands without latency and another program on the control unit received the commands and printed the receiving time to the serial port.
+The command processing time was the difference between the receiving time for each command.
+We sent $1,000$ BLE commands with five UART commands in each of them, and computed an average processing time of $2.96\pm 1.20$ ms.
+Thus, during actual operation of the toolkit, the BLE command sending delay was set to $5$ ms with some tolerance.
+The control unit could simultaneously update five vibration units at $200$ Hz refresh rate, satisfying the signal segmentation algorithm requirements.
 Although vibration units were not directly driven by continuous audio signals from the central MCU, the high bandwidth ensured that they could approximate continuous waveforms using discrete commands (DR2).
 4.3. Power Consumption and Battery Life
-Power consumption measurements would provide an estimation of how long wearable devices built using VibraForge would function before recharging was necessary (DR6).
+Power consumption measurements provide an estimation of how long wearable devices would function before recharging was needed (DR5).
 Power consumption was measured using a bench power supply with a real-time measurement of voltage, current, and power.
@@
 Therefore, for a vibrotactile arm display with one control unit and two full chains of vibration units ($16x2)$, the idle current would be $186$ mA and a $500$ mAh LiPo battery would last $2.69$ hours.
-Even in extreme situations such as multimedia entertainment, running two actuators consistently would consume $486$ mA, leading to $1.03$ hours of operating time.
-For a larger system with four chains of vibration units ($416x$ 4), the idle current would be $266$ mA, so a $8200$ mAh LiPo battery would last $30.83$ hours.
-In an extreme situation where eight actuators were continuously activated, the overall current would be $1460$ mA and the battery would last $5.62$ hours.
+Even in extreme use cases such as multimedia entertainment, continually running two actuators would consume $486$ mA, leading to $1.03$ hours of operating time.
+For a larger system with four chains of vibration units ($416x$ 4), the idle current would be $266$ mA, with a $8200$ mAh LiPo battery lasting $30.83$ hours.
+In an extreme situation with eight actuators continuously activated, the overall current would be $1460$ mA and the battery would last $5.62$ hours.
 4.4. Maximum Number of Vibration Units On Each Chain
 Here we described how the maximum number of units on each chain was determined (DR1).
-Due to incremental wire resistance along the chains, voltage drops will become more significant when additional vibration units are connected or when more units are activated simultaneously.
-To maintain stable operation, the MCU power line must sustain a minimum of $2.3$ V (i.e., the MCU’s minimum operational voltage) and the actuator power line should not fall below $0.9$ V (i.e., the LRA-X actuator’s rated voltage).
-Figure 11. Voltage drops on the MCU power line and the actuator power line with an increasing number of actuators being activated.
+Due to incremental wire resistance along the chains, voltage drops become more significant when additional vibration units are connected or activated simultaneously.
+To maintain stable operation, the MCU supply must maintain above $2.3$ V (i.e., the MCU’s minimum operational voltage) and the actuator supply should not fall below $0.9$ V (i.e., LRA-X actuator’s rated voltage).
+Figure 11. Results of the voltage drops on the MCU (red) and the actuator (blue) power supply with an increasing number of activated actuators.
 […]
 This graph shows the voltage drop along the MCU and actuator power lines as the number of activated actuators increases, highlighting how the voltage decreases significantly for the actuators and slightly for the MCU, with critical thresholds marked for the MCU’s minimum operating voltage and the actuator’s rated voltage.
-We measured the voltage drop threshold by incrementally activating vibration units on the chain.
-An oscilloscope was used to measure the MCU power line and actuator power line at the last vibration unit of the chain.
-If the last unit maintained a stable voltage, then one could be certain that other units were stable as well.
-Figure 11 shows the relationship between voltage drop number of actuators activated.
-When the number of actuators surpassed 16, the MCU voltage dropped below the operating range.
+We measured the voltage drop threshold by incrementally activating vibration units on a chain.
+An oscilloscope measured the MCU power line and actuator power line at the last vibration unit of the chain.
+If the last unit maintained a stable voltage, then one could assume the other units were stable as well.
+When the number of actuators surpassed 16, the MCU voltage dropped below the operating range (Figure 11).
 When it surpassed 17, the actuator voltage dropped below its rated voltage.
@@
 To demonstrate that the toolkit can provide precise control over different vibration characteristics (DR2), we measured the acceleration of vibration units at various intensities and frequencies.
-For intensity, we used an ADXL345 accelerometer for the measurements.
-Since most ADXL345 breakout boards are larger then the vibration unit itself, we designed a FlexPCB version and attached it between the skin and the vibration unit for precise measurements.
+For intensity, we used designed a FlexPCB version of the ADXL345 accelerometer and attach it between the skin and the vibration unit.
 A vibration unit was mounted on a sleeve and placed near the wrist of a human subject.
-During the testing, the vibration unit was activated for three seconds at each intensity level and the vibration frequency was kept constant at 170 Hz.
+During testing, the vibration unit was activated for three seconds at each intensity level and the vibration frequency was kept constant at 170 Hz.
 The average acceleration was measured as the RMS value over the 1-second period of stable vibration.
-Acceleration was found to increase Monotonically as the intensity level increased (Figure 12 Left).
+Acceleration was found to increase monotonically as intensity increased (Figure 12 Left).
+Similarly, for frequency, the vibration unit was placed near the wrist and activated for three seconds at each frequency level with an intensity of 25% .
+The acceleration data was processed using a fast Fourier transform and the resulting main frequency responses were found well aligned with the designated frequency at each level, thereby demonstrating the precise PWM signal control (Figure 12 Right).
+4.6. Waveform Approximation
+Because vibration units are controlled via discrete UART commands rather than continuous audio signals, one might worry that this would reduce the expressivity of the vibration waveform delivered through the units.
+By utilizing high-bandwidth communication (DR2), VibraForge was capable of approximating audio signals and creating similar feedback.
+To demonstrate this, we used standard vibration waveforms from VibViz (Seifi et al., 2015), converted them from WAV files to HAPTIC files using Meta Haptics Studio (Meta Platforms, 2023a), and sent them to the GUI Editor for conversion into a sequence of vibration commands to drive a vibration unit.
+After measuring the acceleration of the vibration unit and comparing it to the original audio envelope, both measurements closely approximated each other (Figure 13).
 Figure 12. The acceleration measurements at various intensity and frequency levels.
@@
 This figure presents two measurements: on the left, a bar chart showing the increase in RMS acceleration as the intensity level rises from 1 to 15, and on the right, a frequency spectrum displaying the magnitude of accelerations at eight different frequency conditions ranging from 125 Hz to 384 Hz.
-Similarly, for frequency, the vibration unit was placed near the wrist and activated for three seconds at each frequency level with level-3 intensity (25% duty cycle).
-To obtain the main frequency response, the acceleration data was processed using a fast Fourier transform.
-We found that the main frequency response was well aligned with the designated frequency at each level, thereby demonstrating the precise PWM signal control (Figure 12 Right).
-4.6. Waveform Approximation
-Because vibration units are controlled via discrete UART commands rather than continuous audio signals, one might worry that this would reduce the expressivity of the vibration waveform delivered through the units.
-Here we show that by utilizing high-bandwidth communication (DR2), VibraForge was capable of approximating audio signals and creating similar feedback.
-We used standard vibration waveforms from VibViz (Seifi et al., 2015), converted them from WAV files to HAPTIC files through Meta Haptics Studio (Meta Platforms, 2023a), and sent them to the GUI Editor.
-The editor automatically converted the haptic files to a sequence of vibration commands that were used to drive a vibration unit.
-We then measured the acceleration of the vibration unit and compared it with the original audio envelope.
-The results showed that they were closely approximated to each other (Figure 13).
-Figure 13. Comparison of actuator vibrations (green lines) and original audio waveforms (light green shaded areas) from the VibViz Library
+Figure 13. Comparison of actuator vibrations (blue lines) and original audio waveforms (light blue lines) from the VibViz Library.
 […]
-This figure compares actuator vibrations (green lines) and original audio waveforms (light green shaded areas) from the VibViz Library, showing four different waveforms.
+This figure compares actuator vibrations (blue lines) and original audio waveforms (light blue shaded areas) from the VibViz Library, showing four different waveforms.
+4.7. Cost Estimation
+VibraForge is a cost-effective toolkit.
+The small control unit costs \$27.23 USD (i.e., \$12.50 for the MCU, \$10.50 for the PCB, \$4.00 for the LiPo batteries, and \$0.23 for the 3D printed parts), which is similar to Arduino Uno but cheaper than Arduino Mega2560 and Raspberry Pi.
+The vibration unit costs \$3.46 USD (i.e., \$2.43 for the PCB, \$1.00 for the LRA, and \$0.03 for the 3D printed parts).
+For the wearable vibrotactile device in Case Study 1, for example, the total cost of a small control unit and 24 vibration units would be \$ 110.27 USD.
+This low manufacturing cost would enable designers to reproduce our toolkit with low financial burden.
+Note that these estimates are based on mass production (e.g., 10 control units and 100 vibration units), so prices may increase when manufacturing in smaller quantities.
 5. Case Studies
-We conducted three case studies to demonstrate the potential usage of VibraForge.
-The first case study showcased how VibraForge can be used to efficiently reproduce previous research prototypes.
-The second and third case studies showed how VibraForge can be used to build novel spatial vibrotactile systems that improve user performance in various applications.
+We conducted three case studies to investigate the potential “ceiling” (Ledo et al., 2018; Myers et al., 2000) of the toolkit’s design space, measured by scalability and expressivity.
+The first case study demonstrated the toolkit’s capability for high expressivity by replicating a phonemic display that requires complex vibration patterns.
+The second case study showcased high scalability by utilizing a larger number of units for VR interaction.
+The third case study found a balance between the two metrics, where parameterized vibrations were used to convey collision risks during drone teleoperation.
 5.1. Case Study 1: Replicating A Phonemic-Based Tactile Display
-In Reed et al.’s work, a phonemic-based tactile display was designed to aid speech communication (Reed et al., 2018).
-This tactile display used 24 actuators on the forearm to deliver unique tactile codes of 39 English phonemes to deaf and hard-of-hearing users.
+In Reed et al.’s work, a phonemic-based tactile display used 24 actuators on the forearm to deliver tactile codes of 39 English phonemes to deaf and hard-of-hearing users(Reed et al., 2018).
 Phonemes were mapped to vibration patterns using frequency, waveform, amplitude, spatial location, and movement.
-Although the display was shown to be useful, the process of creating the display was challenging.
-To drive the 24 actuators independently, the researchers had to use a desktop audio device (i.e., a MOTU 24Ao) with three custom amplifier boards to provide output signals, which made the whole system bulky and unable to be used in real-world situations.
-In this section, we demonstrated how the VibraForge toolkit could be utilized to simplify the development process and enhance the portability of the display.
-5.1.1. Vibrotactile System Development
-The first step in replicating the phonemic tactile display was to determine the layout of the actuators and the chain connections.
-Since the original display used a 4x6 layout on the forearm, we decided to use four chains with six vibration units on each chain and a small control unit.
+Although the display was found to be useful, the building process was challenging.
+A desktop audio device (MOTU 24Ao) with three custom amplifier boards was needed to drive the 24 actuators independently, which made the system bulky and unusable in real-world situations.
+In this section, we demonstrated how the VibraForge toolkit could simplify the development process, enhancing usability and portability.
 Figure 14. The development process of Case Study 1’s phonemic tactile display re-implementation using VibraForge.
 […]
-The top row shows all parts used for assembly, the assembly process itself, and the final prototype worn on an arm.
-The bottom row shows screenshots of waveform creation and layout design for the display, as well as an image of the original prototype used for comparison.
-The second step was to assemble the display.
-The display required a sleeve garment and multiple modules from the toolkit (Figure 14).
-We first assembled the vibration units, then attached them to the garment and connected JST wires between them.
-The control unit was attached to the garment and four chains of vibration units were then connected to the control unit.
-The mapping between chains and connection ports followed the order of actuators described in the original paper.
-The whole assembly process took 16 minutes.
-Once the hardware was built, the next step was to design the vibration patterns.
-Each phoneme needed to represented by a distinctive waveform that incorporated a high-frequency primary component and a low-frequency envelope, with a duration ranging from 100 to 480 ms.
-Consonants were designed as static patterns with multiple actuators being activated concurrently, whereas vowels were designed as dynamic patterns with actuators being activated in succession.
-The GUI Editor streamlined the process of creating the large variety of waveforms.
-For example, to design the waveform for the consonant ”V” (IPA symbol /v/), we first dragged a sine wave into the Waveform Visualizer and set the frequency to be 300 Hz (primary component).
-Then we multiplied it with another sine wave at 8 Hz (envelope).
-The combined waveform was then saved in the Waveform Library as ”Consonant_B”.
-Alternatively, when the envelope was not available in the Waveform Library (e.g., $cos^{2}$ for consonant ”H”), we designed the waveform in Syntacts and imported the .CSV file into our GUI Editor.
-We then used the ”create new chain” function to create a virtual layout with four chains of vibration units on the canvas.
-We selected each of the designated vibration units (i.e., i6, ii6, iii6, iv6), assigned the custom waveform, and set the time range from 0 to 400 ms.
-Finally, we clicked the ”play” button to validate the design.
-The average time for an experienced haptic researcher on our research team to create one phoneme was one minute and five seconds.
-Therefore, the whole phoneme design process could be finished in 42 minutes.
-While the original paper did not mention the waveform design time, we believe the streamlined waveform creation process led to a significant reduction of design time.
+The top row shows the parts used for assembly, the assembly process itself, and the final prototype on an arm.
+The bottom row shows screenshots of waveform and layout creation, as well as an image of the original prototype for comparison.
+5.1.1. Vibrotactile System Development
+The first step was to determine the layout of the actuators and the chain connections.
+Since the original display used a 4x6 layout on the forearm, we decided to use four chains with six vibration units on each chain and a small control unit.
+Then we assembled the vibration units, attached them to the garment, and used JST wires to connect them (Figure 14).
+The control unit was then attached to the garment and four chains of vibration units were connected to the control unit.
+The mapping between chains and connection ports used the order of actuators described in the original paper.
+The assembly process took 16 minutes.
+Once the hardware was built, we designed vibration patterns in the GUI Editor.
+Each phoneme was represented by a distinctive waveform that incorporated a high-frequency primary component and a low-frequency envelope, with a duration between 100 and 480 ms.
+Consonants were designed as static patterns with multiple actuators activated concurrently, whereas vowels were designed as dynamic patterns with actuators activated in succession.
+To design the waveform for the consonant ”V” (IPA symbol /v/), for example, we began by dragging a 300 Hz sine wave into the Waveform Visualizer and multiplying it with an 8 Hz sine wave, saving the result as a custom waveform in the library.
+If an envelope waveform like $cos^{2}$ (for consonant ”H”) was unavailable, we created it in Syntacts (Pezent et al., 2020) and imported it into the Editor.
+We then used the ”create new chain” function to design a layout with four chains of vibration units on the canvas.
+We selected each of the designated units (i.e., i6, ii6, iii6, iv6), assigned the custom waveform, set the duration as 400 ms.
+Finally, we validated the design by clicking the ”Play” button.
+An experienced haptic researcher averaged one minute and five seconds to design a phoneme, and would thus complete the entire process in 42 minutes.
+This streamlined design process would foster rapid design iterations.
 5.1.2. Lessons Learned
-The use of VibraForge to create the phonemic tactile display had several benefits.
-First, compared with the original device, our system was fully portable (DR5), with a small control unit and 24 vibration units directly attached to the sleeve.
-Second, VibraForge simplified the overall development process.
-The hardware system was reduced to module assembly without soldering, and the waveform design was supported by the GUI Editor (DR3).
-Third, the flexible design (DR4) enabled personalization opportunities.
-The original paper mentioned that the inter-motor spacing of the actuators needed to be adjusted for different participants and that the process of adjusting the layout was rather tedious.
-With our system, because the vibration units could be easily attached and detached, researchers could easily adjust the positions even if the garment was worn on the arm, thus giving them much more flexibility.
-Meanwhile, some challenges were found when using the toolkit.
-While designers could easily author patterns using the GUI Editor, it mainly served as a testing site, because the current design of GUI Editor only allows users to manually click the ”play/stop” button to control the vibration sequences.
-This has constrained its use for real world interaction, where speech is segmented into phonemes and used to trigger corresponding patterns in real time.
-Right now the Python server does not support this kind of interaction so an additional API would be needed.
+The use of VibraForge to create the phonemic tactile display offered several advantages.
+First, it was fully portable (DR5) with a compact control unit and 24 vibration units integrated into the sleeve.
+Second, VibraForge streamlined the development process with solder-free module assembly and rapid waveform design in the GUI Editor (DR3).
+Third, its flexible design (DR4) enable easy position adjustments, addressing the design challenge of personalization mentioned in the original paper.
+Meanwhile, some challenges remained.
+While the GUI Editor allowed for easy pattern creation, it mainly served as a testing site and lacked integration into real applications.
+If the designer wanted to play vibration patterns when certain phonemes are detected in speeches, the designer needed to manually convert patterns into sequences of vibration commands to fit the APIs.
+New API functions are needed to support the integration of GUI-designed waveforms into external programming environment.
 5.2. Case Study 2: Fitness Gaming in Virtual Reality
-Virtual Reality (VR) fitness applications like OhShape (Lab, 2019) and Supernatural (Supernatural, 2024) have gained popularity by turning traditional fitness routines into interactive, gamified experiences.
-However, users often face challenges due to the absence of real-time, precise feedback about the physical positioning of their body (Wu et al., 2023).
-While prior research has demonstrated the utility of vibrotactile feedback for movement guidance in rehabilitation applications (Bark et al., 2014; Prabhu et al., 2020), systems used in these applications only deployed actuators on major joint locations such as the wrists or elbows, thus offering limited and coarse guidance.
-Inspired by OhShape (Lab, 2019) and the TV show ”Hole in the Wall” (Fox, 2008), we built a VR fitness application in which participants were asked to adjust their body poses to mimic cutouts in virtual walls when their digital avatar moved through the wall (Figure 15).
-We used the VibraForge toolkit to construct an upper-body suit that would provide spatial vibrotactile feedback to improve training performance.
+Virtual Reality (VR) fitness applications like OhShape (Lab, 2019) and Supernatural (Supernatural, 2024) have turned traditional fitness routines into interactive, gamified experiences.
+However, the absence of real-time and precise feedback about the physical positioning of one’s body makes them difficult to use (Wu et al., 2023).
+While prior work has demonstrated the utility of vibrotactile feedback for movement guidance (Bark et al., 2014; Prabhu et al., 2020; Jo et al., 2023), actuators were only deployed on major joint locations (e.g., wrists or elbows), thus offering limited and coarse guidance.
+Inspired by the TV show ”Hole in the Wall” (Fox, 2008), we built a VR fitness application that asked participants to adjust their body to mimic cutouts in virtual walls (Figure 15).
+VibraForge was used to construct an upper-body suit that provided localized vibrotactile feedback to improve user performance.
 Figure 15.
-An overview of the pipeline used in Case Study 2’s VR fitness game.
-Participant body movements were tracked by Vicon Motion Cameras and sent to a Unity scene for collision detection.
+An overview of the pipeline used in Case Study 2, where participant body movements were tracked by Vicon Motion Cameras and sent to a Unity scene for collision detection.
 Collision commands were then sent to a control unit on the participant’s back through Bluetooth and converted to UART commands for the vibration units.
 […]
-The Vicon motion capture cameras captured one’s motion and transmitted the body tracking data to a Unity scene, where it detected if a collision occurred between the participant and a virtual wall.
-If a collision occurred, then it activated the corresponding actuator to deliver tactile feedback.
+The Vicon motion capture cameras captured one’s motion and transmitted the body tracking data to a Unity scene, where it detected if a collision occurred between the participant and a virtual wall and activated the corresponding actuator to deliver tactile feedback.
 5.2.1. Vibrotactile System Development
-The first step of our development was to determine the layout of vibration units on the body.
-We referenced prior work that measured spatial acuity (Huang et al., 2024a) and determined that vibration units should be uniformly distributed on the arm with 4 cm of spacing between them and on the body with 8 cm of spacing between them.
-As a result, the suit had 36 vibration units on the body and 40 vibration units on each arm (Figure 15).
-The second step was to mount the control units and vibration units on the garment, which was an exercise compression shirt.
-Since the total number of vibration units was within the upper limit of the large control unit, we only needed to use one control unit.
-It was worn on the upper back to help users maintain their balance during fitness training.
-Vibration units were mounted one-by-one and JST wires were used to connect them.
-The third step was to integrate the hardware into the fitness game, which was built using the Unity Game Engine (Technologies, 2020).
-The physical layout of the vibration units on the suit was manually converted into a virtual layout and overlaid on the user’s avatar, so that the virtual vibration units would move following the human body movement.
-Whenever a collision was detected between a virtual vibration unit and an obstacle, the system would determine which vibration units should be activated and send the vibration commands to the control unit through a Python server (Figure 15).
-Since collisions were binary events (i.e., collided / not collided), the vibration intensity was set to constant level (i.e., 7), as well as frequency (i.e., 170 Hz).
+The first step was to determine the suit layout and assemble the hardware components.
+We referenced prior work on spatial acuity (Elsayed et al., 2020; Huang et al., 2024a) and determined that vibration units should be uniformly distributed on the arm with a spacing of 4 cm and on the body with an 8 cm spacing.
+The suit thus had 36 vibration units on the body and 40 on each arm (Figure 15).
+The large control unit was used to drive the substantial number of vibration units, and a compression shirt was used as the garment.
+We affixed the control unit to the upper back to help maintain balance.
+Then we affixed the vibration units and connected them with JST wires.
+Once the suit was assembled, we integrated the hardware into the fitness game, which was built using the Unity Engine (Technologies, 2020).
+The physical layout of the vibration units was manually converted into a virtual layout and overlaid on an avatar so that the virtual units would follow human body movement.
+Whenever a collision was detected between a virtual unit and a wall, the system sent vibration commands to the control unit via Unity plugin APIs and activated corresponding units (Figure 15).
+Since collisions were binary (i.e., collided / not collided), the vibration intensity and frequency were set constant (i.e., 50% and 170 Hz).
 5.2.2. User Study Design
-A preliminary study was conducted to evaluate the effectiveness of the vibrotactile suit on fitness training performance.
-The VR application was run on an Oculus Quest 2 headset (Meta Platforms, 2023b) through a Quest Link connection to a PC (i.e., Alienware M15 Laptop with RTX 3060Ti GPU).
-Vicon Motion Capture cameras (Vicon, 2023) were used to track the body poses of each participant and the poses were then mapped to the virtual avatar in the Unity scene.
-Nine participants (i.e., 4 Males and 5 Females; mean age = 25 years, SD = 3 years) were recruited to participate in the study, which lasted approximately one and a half hours, and each received \$40 compensation.
+A preliminary study evaluated the effectiveness of the vibrotactile suit on fitness training performance.
+The VR application was run on an Oculus Quest 2 headset (Meta Platforms, 2023b) through a Quest Link connection to an Alienware M15 Laptop.
+Vicon Motion Capture cameras (Vicon, 2023) tracked the body pose of each participant and mapped it to a virtual avatar in the Unity scene.
+Nine participants (i.e., 4 Males and 5 Females; mean age = 25 years, SD = 3 years) were recruited for the 1.5-hour study, and each received \$40 compensation.
 The study was approved by our institutional research ethics board and conducted in a university lab space.
 During the study, each participant would experience two conditions, i.e., no feedback and vibrotactile feedback.
-In the no feedback condition, participants only saw the incoming virtual walls and their own avatar and had to rely on proprioception to adjust their body position.
+In the no feedback condition, participants only saw the incoming virtual walls and their own avatar and relied on proprioception to adjust their body position.
 In the vibrotactile feedback condition, participants would feel vibrations at the body positions where collisions occurred between vibration units and walls.
@@
 5.2.4. Lessons Learned
-The benefits of using VibraForge to build this high-density vibrotactile suit were multi-fold.
-First, scalability (DR1) enabled us to support a large number of actuators and provide highly localized feedback to guide body movement.
-Second, the modular and flexible design (DR4) made the system assembly process as simple as press-fitting the vibration units onto the garment and connecting wires between them.
-The entire process took less than 30 minutes.
-Compared with prior work that used sewing techniques (West et al., 2019; Lemmens et al., 2009), our solution was more flexible and efficient.
-Third, the low latency and high bandwidth communication (DR2) helped deliver almost real-time vibrotactile feedback to users when collisions occurred.
-Participants did not report any perceivable delays during the study.
-Nonetheless, there were several challenges during the process.
-First, the integration of the vibrotactile system into the interactive game design was difficult.
-This was because we had to manually convert the physical layout of units on the body to a virtual layout in the Unity Engine and assign vibration units to their positions.
-This process would have been easier if we could detect the physical positions of the units and generate a virtual mapping automatically.
-Second, two participants reported that vibration units near the waist were not consistently perceived because body movement during fitness training would loosen the contact with garment.
-Our temporary solution was to tighten the suit so that movement was constrained.
-In the future, other methods would be needed to ensure that the vibration units maintained contact with the skin.
-5.3. Case Study 3: Teleoperation Assistance for Unmanned Aerial Vehicle Operation
-In this case study, we used VibraForge to design a wearable vibrotactile device to provide Unmanned Aerial Vehicle (UAV) operators with spatial information about obstacles.
-Current UAV teleoperation is dependent on visual feedback from onboard cameras, which presents significant challenges as a camera’s limited field of view can hinder situation awareness and lead to potential collisions with obstacles.
-While there have been prior attempts to use haptic feedback to notify operators of nearby obstacles (Brandt, 2009; Hou et al., 2017; Zhang et al., 2020; Xia et al., 2022), they required an operator to switch from using radio control (RC) controllers to hand-held haptic joysticks, only rendered force feedback in one direction at a time, or had fixed actuator layouts that could not represent lateral movements due to the absence of actuators on the side.
-By creating a new vibrotactile feedback system using VibraForge, one could flexibly attach multiple vibration units to an operator’s body that would render simultaneous spatial cues about surrounding obstacles while the operator was using an RC controller.
-This would enhance operator’s collision avoidance ability and situational awareness in environments where visual feedback is limited.
+Using VibraFogre to build the high-density vibrotactile suit offered several advantages.
+First, scalability (DR1) supported numerous vibration units (116 in total), providing localized feedback for guiding body movements.
+Second, the modular and flexible design (DR4) reduced assembly to less than 30 minutes with simple press-fit attachment and rapid adjustments.
+Third, low latency and high bandwidth communication (DR2) ensured real-time feedback when collisions happened, with no perceivable delays reported.
+Nonetheless, some challenges arose.
+The process of manually mapping the vibration units to the virtual layout in Unity was rather tedious.
+This could be solved by a localization pipeline that automatically detects vibration units’ positions and their relative orders.
+Additionally, two participants reported inconsistent vibrations near the waist due to loosened fabric contact and body movement.
+We addressed the problem by temporarily tightening the suit, but future solutions are needed to ensure consistent skin contact.
+5.3. Case Study 3: Collision Avoidance for Unmanned Aerial Vehicle (UAV) Teleoperation
+In UAV teleoperation tasks, operators heavily rely on visual feedback to detect obstacles and avoid collisions, but the narrow field of view from onboard cameras limits their situational awareness and increases collision risks.
+While prior work attempted to use haptic feedback for notifying operators of obstacles (Brandt and Colton, 2010; Hou et al., 2017; Zhang et al., 2020), they required operators to switch from radio control (RC) controllers to unfamiliar haptic joysticks, and provided only one directional force feedback.
+A spatialized vibrotactile feedback system using VibraForge could deliver simultaneous spatial cues about multiple obstacles, while operators continued using RC controllers (Figure 16).
+Additionally, vibration characteristics can be used to convey different levels of risks, helping operators prioritize high-risk obstacles.
+Figure 16.
+The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point bodily vibrotactile feedback.
+Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange).
+[…]
+A human operator wearing the vibrotactile device is flying a simulated drone.
+He perceives incoming obstacles through vibrotactile feedback triggered on the body.
+He not only sees visible obstacles but also perceives obstacles outside the field of view.
 5.3.1. Vibrotactile System Development
-The first step of the development was to determine the layout of the vibration units on the body.
-Since there is no prior work that studied the mapping between tactile feedback positions on the body and human perception of 3D directions, we conducted a perceptual study with 10 participants to identify a mapping.
+Our first step was to determine the layout of the vibration units on the body.
+Since there is no prior work on mappings between bodily tactile feedback positions and human perceptions of 3D directions, we conducted a perceptual study with 10 participants to identify a mapping.
 The layout was initialized with 46 vibration units uniformly distributed on the body and optimized using user data.
-The final layout used 32 vibration units to achieve comprehensive coverage of the 3D space (Figure 16(b)).
+The final layout used 32 vibration units to achieve comprehensive coverage of the 3D space (Figure 17).
 More details on the perceptual study and layout optimization can be found in (Huang et al., 2024b).
+Then we affixed press-fit 32 VCA vibration units to the garment, which was a compression shirt similar to Case Study 2.
+The vibration units were grouped into four chains, covering the front left, front right, back left, and back right.
+A large control unit was placed on the operator’s back.
 Figure 17.
-(a) The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point vibrotactile feedback on the body.
-Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange).
-(b) The layout of vibration units on the upper body.
+The layout of vibration units on the upper body.
 Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
 […]
-(a) The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point vibrotactile feedback on the body.
-Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange).
-(b) The layout of vibration units on the upper body.
-Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
-The second step was to affix control units and vibration units on the garment, which was a compression shirt similar to that used in Case Study 2.
-Thirty-two VCA vibration units were press-fit to the garment.
-A large control unit was placed on the operator’s back.
-The vibration units were grouped into four chains, each covering the front left, front right, back left, and back right.
-The third step was to integrate the hardware into the simulated environment, which was built using Microsoft AirSim (Shah et al., 2017) with Unreal Engine 4 and Python APIs.
-Since the mapping between vibration units and 3D directions were known from the perceptual study, when a potential collision was detected, the system would trigger the unit that had the nearest direction with the colliding object in the space.
-This avoided the need to reproduce the physical layout of the vibration units in the system.
-The risks of colliding with surrounding obstacles were computed using control barrier functions (Ames et al., 2019) and scaled to $[0,15]$ to match the actuators’ intensity levels.
-If the risk of an obstacle was above $0.5$, the vibration unit in the corresponding direction would be activated by sending commands through a Python server.
+The layout of vibration units on the upper body. The initial layout was designed using 46 units with uniform spacing, while the final layout only used 32 units with more sparse spacing.
+We then integrated the hardware into a Microsoft AirSim (Shah et al., 2018) simulation environment, built with Unreal Engine 4 and our Python server.
+Since the mapping between vibration units and 3D directions were known, when a potential collision was detected, the system would trigger the unit that had the nearest direction with the colliding object in the space.
+The collision risks with obstacles were computed using control barrier functions (Ames et al., 2019) and scaled to $[0,15]$ to match the actuators’ intensity levels.
+If the risk of an obstacle was above $0.5$, the vibration unit in the corresponding direction would be activated by sending parameterized vibration commands through the Python server.
+Figure 18. A first-person view of the simulation environment used in Case Study 3’s study, with randomly positioned obstacles and a top-down view on the right.
+[…]
+This figure shows a navigation scene demonstrating the obstacle course where participants navigate using different control conditions.
+Figure 19.
+The results for the total distance traveled, number of collisions, and input disagreement metrics for the feedback conditions (NA, FSC, VSC) and flying directions (forward-FWD, right-R, and upward-UP).
+The error bars represent the standard error of the mean (SEM), $\mathchoice{{}^{\displaystyle*}}{{}^{\textstyle*}}{{}^{\scriptstyle*}}{{}^{\scriptscriptstyle*}}=p<0.05$, $\mathchoice{{}^{\displaystyle**}}{{}^{\textstyle**}}{{}^{\scriptstyle**}}{{}^{\scriptscriptstyle**}}=p<0.01$, and $\mathchoice{{}^{\displaystyle*{*}*}}{{}^{\textstyle*{*}*}}{{}^{\scriptstyle*{*}*}}{{}^{\scriptscriptstyle*{*}*}}=p<0.001$.
+[…]
+Bar graphs display the performance metrics, including total distance, number of collisions, and input disagreement across multiple conditions (NA, FSC, VSC, etc.), highlighting significant differences between the control methods.
 5.3.2. User Study Design
@@
 For the FSC condition, a Novint Falcon haptic joystick (Inc, 2024) was used for both input and output, similar to previous studies (Omari et al., 2013; Reyes et al., 2015).
-For each condition, the experimental scene was a $5m\times 5m\times 50m$ tunnel (Figure 17(a)) that required participants to steer quadrotors forward, right, or upward.
+For each condition, the experimental scene was a $5m\times 5m\times 50m$ tunnel (Figure 18) that required participants to steer quadrotors forward, right, or upward.
 Different flying directions were used to vary visual information capacity as obstacles are often outside the camera’s field of view when flying right and upward.
-The tunnel contained fifteen objects that were randomly placed to avoid learning effects.
+Fifteen objects were randomly placed in the tunnel to avoid learning effects.
 Based on these conditions, we hypothesize that (a) vibrotactile feedback would enhance teleoperation performance compared to no feedback, and (b) the improvement would be more effective in the right and upward directions than the forward direction.
-Figure 17.
-(a) A first-person view of the simulation environment used in Case Study 3’s study, with randomly positioned obstacles and a top-down view on the right.
-(b) The results for the total distance travelled, number of collisions, and input disagreement metrics for the feedback conditions (NA, FSC, VSC) and flying directions (forward-FWD, right-R and upward-UP).
-The error bars represent the standard error of the mean (SEM), $\mathchoice{{}^{\displaystyle*}}{{}^{\textstyle*}}{{}^{\scriptstyle*}}{{}^{\scriptscriptstyle*}}=p<0.05$, $\mathchoice{{}^{\displaystyle**}}{{}^{\textstyle**}}{{}^{\scriptstyle**}}{{}^{\scriptscriptstyle**}}=p<0.01$, and $\mathchoice{{}^{\displaystyle*{*}*}}{{}^{\textstyle*{*}*}}{{}^{\scriptstyle*{*}*}}{{}^{\scriptscriptstyle*{*}*}}=p<0.001$.
-[…]
-This figure shows an evaluation of a UAV teleoperation system.
-In (a), a navigation scene demonstrates the obstacle course where participants navigate using different control conditions.
-In (b), bar graphs display the performance metrics, including total distance, number of collisions, and input disagreement across multiple conditions (NA, FSC, VSC, etc.), highlighting significant differences between the control methods.
 5.3.3. Analysis and Results
-We evaluated three metrics: total distance travelled, number of collisions, and input disagreement (i.e., the deviation of user steering commands from the safe input range during flight).
+We evaluated three metrics: total distance traveled, number of collisions, and input disagreement (i.e., the deviation of user steering commands from the safe input range during flight).
 Using a two-way repeated measures ANOVA in SPSS, we assessed the effect of feedback condition and flying direction.
-Post-hoc pairwise comparisons with a Bonferroni correction were conducted for significant results (Figure 17(b)).
-The RM-ANOVA determined that feedback condition had a significant effect on total distance travelled ($F(2,22)=3.585,p<0.05$), as did with flying direction ($F(2,22)=17.868,p<0.001$).
+Post-hoc pairwise comparisons with a Bonferroni correction were conducted for significant results (Figure 19).
+The RM-ANOVA determined that feedback condition had a significant effect on total distance traveled ($F(2,22)=3.585,p<0.05$), as did with flying direction ($F(2,22)=17.868,p<0.001$).
 The interaction between feedback condition and flying direction was also significant ($F(1.957,21.527)=4.916,p<0.01$).
@@
 The pairwise comparison between feedback conditions revealed that there was less disagreement with VSC compared to NA ($p<0.01$)
-In summary, VSC caused fewer collisions and less input disagreement compared to NA, and achieved comparable performance with FSC.
-Additionally, the right and upward directions had more collisions than the forward condition, highlighting the increasing reliance users had on haptic feedback when under limited visual capacity.
+In summary, VSC caused fewer collisions and less input disagreement than NA, and achieved comparable performance with FSC.
+Additionally, the right and upward directions had more collisions than the forward condition, highlighting users’ increasing reliance on haptic feedback when visual capacity was limited.
 Thus, our hypotheses were accepted.
+Figure 20. The usability study environment and examples of the diagrams provided to participants during the study.
+[…]
+The usability study environment, with a laptop, a mouse, a box of hardware components, and a display.
+Three slides used in the study showing diagrams of the design task.
+A task overview with a sketch illustration of the tactile sleeve, a sleeve layout guiding the hardware assembly, and two waveform illustrations for the vibration patterns.
 5.3.4. Lessons Learned
-The process of using VibraForge to build the teleoperation vibrotactile system demonstrated several benefits.
-First, the flexible design (DR4) allowed us to iterate on the layout of vibration units rapidly during the perceptual study.
-From the initial 46 actuator layout to the final 32 actuator layout, we could easily adjust the positions of vibration units on the body and test various spatial mappings.
-Second, the low latency vibration responses (DR2) also provided instant feedback for participants, which is essential in high speed teleoperation tasks.
-In addition, the varied intensities (DR2) provided more nuanced information for participants, helping them prioritize their responses towards obstacles with higher risks.
-Together, these benefits demonstrate the usability of the toolkit for a complex application.
-We also encountered several design challenges.
-One issue was that participants reported that the same intensity level was perceived differently across body areas.
-Locations with bone contacts were perceived to be stronger, while those with more subcutaneous tissues were perceived to be weaker.
-This variability may affect feedback precision during critical applications, and require a deeper investigation into how this can be overcome.
-Another issue was encountered during the layout iteration.
-Because the final layout was relatively sparse, the original connection wires that were used were too short.
-Therefore, we had to solder and assemble new wires.
-This, however, helped us make the toolkit adaptable to more diverse future use cases.
+Using VibraForge to build the teleoperation vibrotactile system revealed several benefits.
+The flexible design (DR4) enabled rapid iterations of vibration unit layouts, from an initial 46-actuator setup to a final 32-actuator configuration, allowing for quick testing of various spatial mappings.
+Low-latency transmission (DR2) provided immediate feedback essential for high-speed tasks, while varied intensities (DR2) offered nuanced information, helping participants prioritize high-risk obstacles.
+These features highlight the toolkit’s capability for complex applications.
+However, design challenges did arise.
+Participants noted that intensity perception varied across body areas, with stronger sensations near bones and weaker ones in areas with more tissue.
+Thus, individual calibrations might be needed to ensure feedback consistency.
+Additionally, the sparse final layout required soldering longer JST wires to replace the original ones.
+This suggested that VibraForge needs more enhancements to support sparse layout designs.
+6. Usability Study
+To understand the usability of the toolkit and determine its ”threshold” (i.e., a user’s ability to get started using the toolkit (Ledo et al., 2018; Myers et al., 2000)), we conducted a first-use lab study (Hartmann et al., 2006) with ten participants recruited from a local technology exhibition.
+Participants included five males and five females with ages ranging from 18 to 40 years old (Mean = 25.6 years, SD = 6.7).
+Participants had a range of prior experience programming, building physical devices, and building haptic/tactile devices, with some never having done so and others proclaiming themselves as experts.
+Each study lasted one hour and each participant received \$30 as compensation.
+The study was approved by our institutional research ethics board.
+6.1. Study Design
+The study consisted of a design task, a free exploration activity, and a post-study questionnaire and interview.
+In the design task, participants replicated a simplified version of the phonemic tactile sleeve display from Case Study 1 (Reed et al., 2018): assembling two chains with six vibration units each and designing two phonemes, consonant /k/ and vowel /au/).
+Participants received a 5-minute tutorial on the toolkit’s hardware components (i.e., control units, vibration units, connector direction, garment attachment, etc.) before assembling the tactile device following a diagram displayed on a computer monitor.
+After assembly, participants underwent another 5-minute tutorial on the GUI Editor and waveform creation workflow.
+They then designed and tested the two phonemes using diagrams on the computer monitor.
+For the 15-minute free exploration activity, participants could assemble new devices on other garments (e.g., headband, glove, socks, etc), or combine multiple waveforms to produce complex patterns in the GUI Editor.
+Participants were encouraged to utilize their creativity, so the experimenter did not provide specific guidance but rather served as an assistant helping participants achieve their design goal.
+The experimenter recorded the creation process and the final outcome from the exploration.
+After that, participants completed a questionnaire with demographic questions and questions from the System Usability Scale (SUS) (Brooke, 2013).
+A semi-structured interview then focused on the overall design experience, key areas for improvement, and potential applications of VibraForge.
+6.2. Results
+All participants successfully completed the design task and free exploration activity.
+The average completion time for the design task’s hardware assembly was 8 minutes 39 seconds (SD = 1 minute 52 seconds) and 8 minutes 3 seconds (SD = 2 minutes 46 seconds) for the design task’s software waveform design.
+These completion times were comparable to the expert time reported in Section 5.1.1.
+Furthermore, an average SUS score of 76.75 was obtained, which is between the 70th and 80th percentile and higher than average SUS scores (Brooke, 2013).
+These results indicate that the toolkit has a very low ”threshold” for participants.
+Overall, participants appreciated the intuitive design of the hardware components and found the assembly process straightforward and intuitive (e.g., P10: ”the hardware was very easy to assemble and it was very intuitive to configure”; P2: ”it feels like LEGO pieces clicking together”).
+Participants also appreciated that the software was easy to learn and demonstrated good interface design (e.g., P1: ”the software part was nice because it was similar to the designing tools like Canva or Figma”; P6: ”the timeline at the bottom was very helpful for me when I was trying to debug what was going on”).
+During the free exploration activity, participants created a variety of haptic devices (Figure 21).
+For example, P4 designed a glove with four vibration units on the dorsal side of the hand.
+P7 attached two vibration units on the left and right sides of the heel on a sock to notify no-entry area.
+P10 designed a glove with three units on the dorsal side and one unit on the palm to ”play” the Queen song ”We Will Rock You” where each note triggered a different unit on the hand.
+P2 used the GUI Editor to transcribe the Nat King Cole song “On a Bicycle Built for Two” and played it using the frequency and duration of notes on each vibration unit.
+P8 made attempts to reproducing other phonemes such as /sh/, /ei/ and /p/ from the original paper (Reed et al., 2018) using the GUI editor.
+Figure 21.
+The devices created by participants during the free exploration: a glove with 4 vibration units on the dorsal side (P4), a sock with two units on the left and right side of the heel (P7), and a glove with three units on the dorsal side and one unit on the palm for transcribing songs (P10).
+[…]
+The devices created by participants during the free exploration activity: a glove with 4 vibration units on the dorsal side (P4), a sock with two units on the left and right side of the heel (P7), and a glove with three units on the dorsal side and one unit on the palm for transcribing songs (P10).
+Participants did note some areas for improvement.
+For the hardware components, participants mentioned that it would be helpful to have more apparent visual feedback on the chain connections (e.g., P2: “I feel like it [chain connector index] would be much more visible if it’s printed on top of it rather than in the middle.; P7: “I didn’t know whether they [vibration units] were connected … I think making [the LED] more visible would be more helpful”.
+They also suggested other methods to affix the vibration units to garments to make the toolkit more adaptable to other fabrics (e.g., P7: ”This [bottom ring attachment] and a magnet would have been a bit more effective, because you would only need to press it … and it’s gonna have a really strong grip on it for you to move it around”; P4: ”maybe there’s like a lock method so, when you snap on it locks into the sensors [vibration units]”.
+Based on these feedback, there is thus a need to add visual indicators for hardware connection statuses and explore additional mechanical affixation structures.
+Others expressed the need for more versatile functionality in the GUI Editor.
+For example, P1 wanted to directly manipulate the waveforms (e.g., ”if I want to change the amplitude, for example, I should just drag the top part and then drag down to lower, like direct manipulation … if I want to change the frequency, I should just click and drag the amount to left or right”.
+Similarly, because participants needed to drag and drop the same waveform onto several vibration units, P6 suggested adding ”multi-select” functions to edit multiple vibration units concurrently.
+Because the Waveform Visualizer was used to both edit and assign waveforms, some participants were confused about its current status.
+P5 commented that ”the fact that the designing of a custom wave and assigning it to the node share the same interface is a bit confusing”.
+These comments suggest that the GUI Editor requires some refinement to ensure that supported actions are more natural and clearly conveyed to users.
+Participants also highlighted several potential applications of the toolkit.
+P1 and P4 noted how VibraForge could be used to aid visually impaired individuals with navigation.
+P2 and P10 demonstrated how the toolkit could be used to translate music into tactile patterns, whereas P6 suggested using it to enhance VR and gaming immersion by simulating action feedback.
+P3 and P9 proposed the creation of devices for personalized massages and stress relief.
+P7 was enthusiastic about using it for safety applications, such as non-visual notifications while running or cycling, particularly in noisy environments.
+These examples demonstrate VibraForge’s potential usage in diverse domains.
 7. Discussion
-In this section, we discussed the implications drawn from case studies, including strengths and limitations of VibraForge.
-Additionally, from our first-hand experiences in case studies, we synthesized a generalized five-step pipeline for building spatialized vibrotactile systems.
-Potential applications of VibraForge were highlighted at the end.
+Herein we discuss the strengths and limitations of VibraForge summarized from the studies, and suggested potential solutions to address the limitations.
+Additionally, based on our first-hand developing experiences, we synthesized a generalized five-step pipeline for building spatialized vibrotactile systems.
 7.1. Strengths of VibraForge
-VibraForge aims to provide a ”least resistance path” (Ledo et al., 2018) towards creating spatialized vibrotactile systems, which was demonstrated through the three case studies.
-VibraForge offers great scalability.
-Systems designed in the case studies are of different scales, ranging from sleeve-type phonemic display with 24 vibration units on the forearm, to body-scale system with over 100 vibration units spanning across the upper body.
-Compared existing toolkits that are often constrained intrinsically to less than 16 actuators, VibraForge’s chain connection method enables designers to easily scale their system to more than 100 vibration units, significantly expanding the potential design space.
-Furthermore, adding a new unit to an existing system is also as simple as plug-and-play, without the need for any soldering or programming.
-VibraForge offers great expressivity.
-Information of various complexity are adequately conveyed through spatial vibrotactile feedback in the case studies, ranging from simple on/off status for VR collision detection, to continuous risk levels in teleoperation and symbolic phoneme representation in the phonemic display.
-Compared with prior spatialized vibrotactile systems that are constrained to using ERMs and low-bandwidth control, our toolkit offers various actuator options and provides fine-grained control over vibration characteristics of each actuator.
-This enables vibration units to reliably reproduce audio waveforms in technical evaluations.
-The GUI Editor of VibraForge features usability.
-In the phonemic display study, it effectively expedited the pattern authoring process by reducing the average pattern authoring time to one minute.
-Compared with prior design interfaces, our editor highlighted the canvas and the timeline that provided substantial support for multi-actuator pattern design.
+VibraForge aims to provide a ”least resistance path” (Ledo et al., 2018) towards creating spatialized vibrotactile systems via bridging the gap between scalability and expressivity.
+The systems designed in case studies ranged from a sleeve-type phonemic display on the forearm, to a body-scale system spanning the upper body.
+Compared to existing toolkits that are often constrained to fewer than 16 actuators (Wittchen et al., 2022; Pezent et al., 2020; Dementyev et al., 2021), VibraForge’s chain connection method enables designers to easily scale their system to more than 100 vibration units, significantly expanding potential design spaces.
+Furthermore, new units can be plug-and-play into existing systems without additional programming, fostering rapid prototyping process.
+VibraForge also features great expressivity.
+It rendered tactile information at different complexities, ranging from simple on/off status for VR collision detection, to parameterized risk levels in teleoperation and symbolic phoneme representation in the phonemic display.
+Compared with previous spatialized vibrotactile systems that were restricted to using ERMs and PWM control (Wang et al., 2021; Sánchez and Bohne, 2015; Lemmens et al., 2009), our toolkit offers various actuator options and provides fine-grained control over vibration characteristics, thus reliably reproducing complex waveforms.
+Furthermore, VibraForge demonstrates high usability.
+Case Study 1 and the usability study indicated that both professional and novice designers benefited from the streamlined hardware assembly and intuitive GUI design.
+Compared with prior design interfaces (Pezent et al., 2020; Minamizawa et al., 2012), our GUI Editor provided substantial support for multi-actuator pattern design.
+Professionals using the GUI Editor effectively reduced the average pattern authoring time.
+Novice users compared the intuitive hardware assembly to the joyful process of ”building LEGOs”.
+Within an hour, they learned the toolkit basics and utilized it to finish professional design tasks and for creative exploration.
+This indicated that the toolkit lowered the technical barrier to building spatialized vibrotactile systems and enhanced the overall efficiency.
 VibraForge also features portability and flexibility.
-Compared with the original phonemic display that was wired to a bulky desktop audio device, components used in our prototype were more portable and easy to wear, such as the small control unit that communicates via Bluetooth.
-In the UAV teleoperation study, flexibility ensured that vibration units could be frequently adjusted during the layout optimization perceptual study.
-Compared to using commercial devices that often have fixed feedback positions, VibraForge enables designers to easily adjust the positions of vibrations units, thus enabling rapid design iterations.
+Compared with the original phonemic display that was wired to a bulky desktop audio device, our components were more portable and easy to wear, such as the small control unit that communicated via Bluetooth.
+In the UAV teleoperation, flexibility ensured that vibration units could be frequently adjusted during the layout optimization perceptual study.
+Compared to using commercial devices (bHaptics, 2024) that often have fixed feedback positions, VibraForge enables designers to easily adjust the positions of vibrations units, thus enabling rapid design iterations.
+In summary, rather than contributing one incremental technical advancement, our work fuses multiple components—modularity, usability, and portability—into a cohesive vibrotactile prototyping toolkit.
+This holistic system-level integration enables a larger design space with a higher “ceiling” (shown in the case studies) and lower “threshold” (shown in the usability study), opening new possibilities for both professional designers and novice users to build applications with haptic feedback.
+VibraForge could be useful within several domains.
+For example, In VR social platforms, spatial vibrotactile feedback could help users express their emotions through social touch interactions (Sykownik and Masuch, 2020).
+Developers using the toolkit can easily adjust the layout of actuators to build personalized devices, in order to avoid discomfort or sensory issues in certain body areas.
+As another example, VibraForge can be used to build prototypes for tactile perceptual studies, such as those measuring spatial acuity, information transfer, or other perceptual characteristics.
+For example, Park et al.
+performed perceptual studies to measure information transfer across five body positions (Park et al., 2023), but the choice of positions were limited by the layout of bHaptics modules (bHaptics, 2024).
+By using VibraForge, researchers could freely adjust the layout of vibration units on the body and adapt it to any study requirements.
 7.2. Limitations of VibraForge
-Although VibraForge can enable designers to quickly prototype and evaluate new multi-actuator vibrotactile devices and experiences, the toolkit does not come without limitations.
-For example, the open-loop control used in the chain connection topology does pose some issues.
-First, when a control unit sends a vibration command, it is unclear if the command was successfully transmitted to the target vibration unit or not.
-This is especially important when sending a ”stop” command, because the vibration unit won’t stop vibrating if the commands was missed during the transmission process.
-Second, when multiple actuators were activated, there was a noticeable voltage drop along the chains.
-Although the technical evaluation showed that the supply voltage remained in the MCU operating range and actuator voltage range, it would still affect the users’ subjective experience of the vibrations.
-By adding extra command receiver pins on the control unit, one could instead use closed-loop control to solve these problems.
-With such a design, the vibration units and the control unit would form a loop, wherein the last vibration unit of the chain would be connected back to the control unit using a receiver pin.
-Whenever a vibration unit receives a command for itself, it would not only process the command and trigger the vibration, but also pass a confirmation message to the next unit.
-When the last unit in the chain receives this confirmation message, it would send the message back to the central unit, thus closing the loop.
-In cases where commands fail during the transmission process, the central unit would miss the confirmation message and resend the message.
+The toolkit does not come without limitations.
+The open-loop control used in the chain connection poses some issues.
+For example, when a control unit sends a vibration command, it can be unclear if the command was successfully transmitted to the target vibration unit.
+When transmitting a ”stop” command, this is especially important as a vibration unit will not stop vibrating if the commands was missed.
+Additionally, when multiple actuators are activated there is a noticeable voltage drop along the chains, which would affect users’ experience of the vibrations.
+By adding extra command receiver pins to the control unit, the vibration and control unit would form a loop, wherein the last vibration unit in a chain would be connected to the control unit using a receiver pin and could propagate a confirmation message back to the central unit.
+In cases where commands fail during transmission, the central unit could resend the message.
 A closed loop connection would also reduce the voltage drop along the chain, as both ends of the chain would be connected to the power source.
-An LTspice simulation shows that an open-loop, eight-unit chain would result in voltage drop from 5 V to 3.1 V, while a closed-loop version would increase the voltage to 4.2 V.
-Additionally, while the GUI Editor’s waveform design was advantageous for the phonemic display pattern design, designers would face difficulties when designing more complex spatial patterns.
-For example, if a designer wanted to create the effect of actuators vibrating from a central point and expanding out in concentric circles, a designer would have to manually design the waveform, assign the waveform to each vibration unit, and specify the start and end time, which is very tedious.
-A better solution would be to use graphical design metaphors and breakdown spatial patterns to low-level animations of point, line, and shapes.
-By leveraging phantom sensation illusions (Schneider et al., 2015), designers would only need to draw the shapes and trajectories, and they would be automatically converted to vibration commands for each unit.
-In this way, complex patterns could be created with ease.
-Our case studies also emphasized the need for an automation pipeline to detect the positions of vibration units on the body and convert them to virtual layout that can be later used in various programming environment.
-We believe this need can be fulfilled with computer vision and motion capture techniques in the future.
-Using RGB images of the body, computer vision methods could detect the human body outline, segment vibration units from the body, and record the units’ positions.
-Similarly, retro-reflective markers could be mounted on top of each vibration unit to create a tracking object in the motion capture system.
-These positions could then be tracked in real-time and loaded into a development environment such as Unity Engine.
-In this way, designers could focus on the interaction design rather than manually creating the virtual layout.
+An LTspice simulation showed that an open-loop, eight-unit chain would result in voltage drop from 5 V to 3.1 V, while a closed-loop would increase the voltage to 4.2 V.
+While the GUI Editor’s waveform design was advantageous for the phonemic display’s pattern design, if a designer wanted to create an effect where actuators vibrated out from a central point in concentric circles, a designer would have to manually assign the waveform and specify the start and end time for each vibration unit.
+Instead, graphical design metaphors could be used to break down spatial patterns into low-level animations that would be automatically converted to vibration commands for each unit.
+VibraForge was built focusing on dense matrices of vibration units, as this was consistent with prior spatial tactile feedback systems that employed dense actuator arrays (e.g., facial masks (Wang et al., 2021) and bodysuits (bHaptics, 2024; Lemmens et al., 2009; West et al., 2019)).
+However, Case Study 3 revealed that it was challenging to support the sparse layout of vibration units on the body, as longer wires were needed to connect more distant units, which may impact the overall system reliability.
+To overcome this limitation, we plan to support the use of multiple control units within the same system so that designers can place control units at various positions on the body to support such sparse layouts.
+Our studies also emphasized the need for an automated pipeline to visualize vibration unit positions.
+Currently, the GUI Editor canvas provides an approximate virtual layout of vibration units and their relative order within each chain.
+Although participants in our usability study were able to map physical units to their virtual counterparts, a true-to-scale, integrated visualization would reduce cognitive load and foster more rapid design iterations.
+One potential solution could be to revise PCB designs and add programmable LEDs to vibration units.
+When units are activated sequentially on each chain, computer vision algorithms could detect the positions of units on the body and their relative orders.
+Alternatively, ultra-wideband (UWB) tags could be mounted on top of the vibration units.
+With multiple UWB anchors setup in the environment, each UWB tag could utilize trilateration to find its own position.
+We are actively exploring these solutions and will incorporate them into the next iteration.
+Lastly, VibraForge currently only supports vibrotactile actuators.
+Recent multimodal haptic toolkits, such as SleeveIO (Shtarbanov et al., 2023) and TactorBots (Zhou et al., 2023), demonstrated the potential of integrating diverse actuator types (e.g., pneumatic, hydraulic, servo-driven) into haptic systems.
+However, these toolkits often rely on direct connections between modules and controllers, which limit their scalability.
+In contrast, VibraForge’s chain-connection architecture naturally supports extensions.
+Adding new actuator modalities would simply require custom PCB drivers and revised data protocols.
+Data bytes for intensity and frequency could be repurposed to handle parameters such as servo speed and position, Peltier temperatures, mechanical displacements, and so on.
+Thus, VibraForge could achieve the expressivity possible with other multimodal toolkits while maintaining its scalability.
 7.3. Pipeline for Creating Spatialized Vibrotactile Systems
-Synthesizing the processes followed during the three case studies, herein we summarize a generalized five-step pipeline for creating spatialized vibrotactile systems.
-The first step is to determine what information to be delivered through vibrotactile feedback.
+Synthesizing the processes followed during the case studies, we summarize a generalized five-step pipeline for creating spatialized vibrotactile systems.
+The first step is to determine what information to deliver using vibrotactile feedback.
 For example, the purpose of the phonemic display was to convey vowels and consonants to deaf and hard-of-hearing people, so the types of vowels and consonants in words were the target information.
-In the drone teleoperation scenario, the key information was the potential risk of colliding with nearby obstacles and could be further divided into the directions of obstacles and their risk levels.
 The second step is to decide how to convey the information.
-Since vibrations can be configured with different characteristics such as intensity, frequency, duration, position, and so on, the common practice is to breakdown the information and deliver it through multiple channels to maximize information transmission (Tan et al., 2020).
+Since vibrations can be configured with different characteristics such as intensity, frequency, duration, position, and so on, the common practice is to break down the information and deliver it via multiple channels to maximize information transmission (Tan et al., 2020).
 For example, vowels and consonants for the phonemic displays were distinguished by using moving patterns and static patterns, while different consonants were distinguished by actuator position, frequency modulation and vibration duration.
-As shown in prior research, providing information through multiple channels with fewer levels in each channel can enhance the overall information transfer for vibrotactile systems.
 The third step is to decide where to install the system and derive the layout of vibration units.
 In cases where the information was spatially connected to the surrounding environment, the layout should have a good mapping between tactile positions and spatial direction.
-For example, in the VR fitness application, vibration units were mainly placed on the arm to provide localized feedback, indicating the arm pose differences between the user and the wall.
 In the UAV teleoperation, locations of vibrations were perceptually mapped to the directions of incoming obstacle.
@@
 The fourth step is to build the system.
-Control units and vibration units can be chosen based on the scale of the system.
-Learning from the building process of case studies, We recommend using the small control unit for localized systems such as sleeves and headbands, as well as applications that require short-term usage.
-Alternatively, the large control unit is suitable for long-term usage and body-scale systems, such as immersive virtual experience and robot teleoperation.
-When designing vibration patterns, if the information involves complex waveforms, designers can utilize the waveform library in the GUI Editor to expedite the authoring process, just like how we designed patterns for the phonemic display.
-If the information only requires changing levels of intensity or frequency, the Python server with APIs should be handy.
+Control and vibration units should be chosen based on the scale of the system (e.g., the small control unit for localized systems such as sleeves and headbands and the large control unit for long-term usage and body-scale systems such as immersive virtual experience and robot teleoperation).
+When designing vibration patterns, if the information involves complex waveforms, designers can utilize the Waveform Library in the GUI Editor to expedite the authoring process, similar to its use for the phonemic display.
+If the information only requires parameterized control, the Python server and APIs should be handy.
 The final step is to test the system and iterate on the design.
-If the desired information cannot be transmitted as expected, the designer can reconfigure the system by changing the layout or experiment different characteristics for conveying the information.
-6.4. Potential Applications for VibraForge
-A toolkit such as VibraForge could benefit several domains.
-In VR, spatial vibrotactile feedback can enhance immersive social experiences.
-In VR social platforms, adding tactile interactions can help users express their emotions through social touch gestures (Sykownik and Masuch, 2020).
-If developers are creating devices for users with sensory issues in certain areas of their body, they can leverage the toolkit to quickly and easily adjust the layout of actuators to avoid discomfort.
-During human-robot interaction, spatial feedback could be useful for assisting humanoid robot teleoperation.
-Humanoid robot training collects data from human operators performing tasks remotely.
-However, the absence of tactile feedback mechanisms makes it difficult for operators to execute daily tasks that are typically performed with ease using their own bodies (Pacchierotti and Prattichizzo, 2024).
-Spatial tactile feedback can provide additional information about remote interaction, such as when a robot hits a wall or other obstacle.
-The choice of tactile feedback positions can also be configured to match the positions of sensors on the robot.
-This feedback could thus increase an operator’s situational awareness and close the interaction loop.
-Last but not least, we envision that VibraForge can be used to build custom prototypes for tactile perceptual studies, such as those measuring spatial acuity, information transfer, or other perceptual characteristics.
-For example, Park et al.
-performed perceptual studies to measure information transfer across five body positions (Park et al., 2023).
-These positions were not chosen by the researchers but instead were constrained by the design of the commercial system they used (i.e., tactile modules from bHaptics (bHaptics, 2024)).
-By using VibraForge, researchers could freely adjust the layout of vibration units on the body and adapt it to any test they want to conduct.
+If the desired information is not transmitted as expected, designers can consider changing the layout or using different vibration characteristics to convey the information.
 8. Conclusion
-Acknowledging the limits with existing spatialized vibrotactile feedback systems and toolkits, we built the VibraForge toolkit.
-Our toolkit enables scalability via a modular design and the use of chain-connection data transmission methods.
-Technical evaluations validated the toolkit design, and case studies demonstrated the wide range of experiences that could be authored.
-The lessons learned during the case studies contributed to our understanding of a generalized pipeline for creating spatialized vibrotactile systems, and highlighted potential research directions of toolkit refinement and future applications.
-The toolkit is in its final stages of preparation, and will be released to the public later this year.
-We are excited to see how researchers and designers will harness the toolkit to elevate their own research and push the boundary of design innovation when using spatialized vibrotactile feedback.
+The VibraForge toolkit enables scalability and expressivity via a modular design and a chain-connection method.
+Technical evaluation validated toolkit’s design requirements, case studies demonstrated the design space, and usability study highlighted low technical barrier and customizability.
+The lessons learned from evaluations contributed to potential directions for toolkit refinement and generalized pipeline for creating spatialized vibrotactile systems.
+We are excited to see how researchers and designers will harness the toolkit to elevate their own research and push the boundaries of design innovation using spatialized vibrotactile feedback.
 Acknowledgements.
+We would like to express our deepest gratitude to Dr.
+Michelle Annett, whose assistance significantly enhanced the quality of our manuscript.
+We would also like to thank Jeb Thomas and Brenna Li for providing feedback on paper writing.
+We also thank Mengfei Liu, who kindly provided the sewing machine for tailoring the garments used in this work.
+Some figures in the manuscript were designed and improved by For Real Entertainment and Annie Squarecircle.
+ChatGPT was used in the writing process for language enhancement.